### v2

# Empowering Scientific Workflows with Federated Agents

Abstract.
Agentic systems, in which diverse agents cooperate to tackle challenging problems, are exploding in popularity in the AI community. However, the agentic frameworks used to build these systems have not previously enabled use with research cyberinfrastructure. Here we introduce Academy, a modular and extensible middleware designed to deploy autonomous agents across the federated research ecosystem, including HPC systems, experimental facilities, and data repositories. To meet the demands of scientific computing, Academy supports asynchronous execution, heterogeneous resources, high-throughput data flows, and dynamic resource availability. It provides abstractions for expressing stateful agents, managing inter-agent coordination, and integrating computation with experimental control. We present microbenchmark results that demonstrate high performance and scalability in HPC environments. To demonstrate the breadth of applications that can be supported by agentic workflow designs, we also present case studies in materials discovery, decentralized learning, and information extraction in which agents are deployed across diverse HPC systems.
Computational Workflows, Distributed Computing, Federated Computing, Multi-Agent Systems, Open-Source Software
1. Introduction
The desire to automate scientific processes has led to advancements in many fields, from artificial intelligence (AI) (Zvyagin et al., 2023) and computational workflows (Deelman et al., 2009) to research data management (Allen et al., 2012) and self-driving laboratories (SDL) (Abolhasani and Kumacheva, 2023), but humans typically remain responsible for core aspects of the iterative research cycle, including hypothesis generation, experimental design, code development, and data analysis. Often, the human-in-the-loop is the rate-limiting step in discovery. This friction increases as the scale and ambition of computational science endeavors grow and leads to inefficient use of research cyberinfrastructure—the federated ecosystem of experimental and observational facilities, data repositories, and high-performance computing (HPC) systems (Miller et al., 2023).
Figure 1. Cooperative agents, spanning federated research infrastructure (experimental facilities, computational systems, data storage), can enable agentic workflows that autonomously steer discovery.
[…]
[Agents deployed across research infrastructure, including experimental facilities, computational systems, and data storage, can facilitate agentic workflows that autonomously engage research infrastructure.]
Intelligent agents, either as an individual system or composing larger multi-agent systems (MAS), rather than humans, can be the driving entities of discovery. Agents are independent, persistent, stateful, and cooperative—working together to achieve a predefined goal with only intermittent human oversight. The contemporaneous explosion of interest in multi-agent systems is largely a consequence of advancements in reasoning capabilities of the large language models (LLMs) often used to back AI agents. Expressing components of scientific applications as agents—programs that can perform tasks independently or semi-autonomously on behalf of a client or another agent—is powerful. An agent manages its own local state and exposes a well-defined behavior. Agents can perform human roles in iterative scientific processes (Wang et al., 2023) or encapsulate research cyberinfrastructure (e.g., computational resources and procedures, experimental instruments, data repositories) (Gao et al., 2024).
Significant progress has been made towards developing AI agents that can act on behalf of humans for such tasks as literature synthesis (Lála et al., 2023), hypothesis generation (Google Research, 2025), and data analysis and publication (Ifargan et al., 2025). However, existing agent frameworks (e.g., AutoGen (Wu et al., 2023)) are not ready to build and deploy agents that employ federated research cyberinfrastructure. New middleware is needed to enable agentic workflows that seamlessly integrate experiment, observation, theory, simulation, AI, analysis, and more, as in Figure 1.
Frameworks for building agentic workflows are limited in scope and generally target conversational, cloud-native applications (e.g., LLM-based AI chatbots) (LangChain, 2024; Wu et al., 2023; OpenAI, 2024). The federated nature of research infrastructure poses unique challenges: distributed resources have diverse access protocols, interactions between computational and experimental entities are asynchronous, and the dynamic availability of resources requires fault-tolerant and adaptive systems. Existing frameworks fail to address these intricacies. They lack abstractions and mechanisms tailored to support autonomous multi-agent workflows that integrate computation, data management, and experimental control, which leads to brittle, ad hoc integrations that are ill-suited for the demands of modern science. Moreover, the inherent complexity of such workflows is compounded by the need to balance efficiency with scientific rigor, especially in applications involving real-time decision-making, iterative exploration, and multi-agent coordination.
These challenges are often orthogonal and span many levels of abstraction, but achieving this vision where intelligent agents serve as driving entities in scientific discovery requires a paradigm shift in how workflows are designed, orchestrated, and executed. We introduce a novel framework for building agentic workflows, emphasizing modularity, statefulness, and interoperability across the diverse research infrastructure. Specifically, this work contributes:
  • Academy, a novel, modular, and extensible middleware for expressing agentic workflows and deploying multi-agent systems across federated resources. Academy addresses unique challenges in scientific applications, such as high data volumes, variable resource availability, and the heterogeneous nature of experimental and computational systems (Section 3).
  • Performance analysis of Academy in diverse scenarios yielding insights into the scalability and practical considerations of deploying agentic workflows (Section 4).
  • Case studies demonstrating the utility of agentic workflow design and highlighting improvements in automation, resource utilization, and discovery acceleration (Section 5).
These contributions advance the state of the art in multi-agent systems for scientific discovery and establish a foundation for future innovations in autonomous research workflows.
2. Background
Agents encompass a rapidly expanding front for AI research, yet agent paradigms can address a breadth of challenges across the computational sciences. We begin with a definition of an agent—inspired by prior work—that is sufficiently generic to encompass the various semantic uses of the term. Then, we enumerate common high-level classes of agents and formalize agentic workflows, both of which we aim to support in the design of Academy in Section 3.
An agent is a program that can perform actions independently or semi-autonomously on behalf of a client or another agent. This definition is imprecise but presents a powerful conceptual model for distributed computing. The agent concept originates from the actor model, a concurrent computing paradigm in which actors encapsulate a local state and communicate through message passing (Hewitt et al., 1973). Agents extend the actor model with the notion of agency—the ability of the agent to engage independently with its environment.
An agent $a$ is defined by its behavior $B$ and local state $S$. The behavior of an agent encompasses a set of actions $x\in X$ (procedures that the agent can perform), and a set of control loops $c\in C$ that define the autonomous behavior (Goodwin, 1995). Agents are often long-running, but may also be ephemeral—created to complete a specific task and then exiting. Clients and agents can request another agent to perform an action on their behalf through message passing. An action can be atomic or composite, invoking other actions on the same or remote agents. An agent with actions but no control loops (i.e., $|X|>0$ and $|C|=0$) reduces to an actor.
Agents come in many flavors. Intelligent (deliberative) agents are goal-oriented and reason about what actions to take using internal models and external perception (Wooldridge and Jennings, 1995). AI agents, a subset of intelligent agents, use AI methods to make decisions or perform actions. In contrast, reactive (observer) agents simply perceive their external environment and react to changes (Nwana, 1996). Service agents provide predefined services in response to action requests and come in many forms: resource agents manage and grant access to resources, such as compute or storage, and embodied agents can act in the world, such as through physical actions when paired with a robot body. Learning agents adapt their behavior over time to improve performance, often through reinforcement learning (Mnih et al., 2015). Composite agents exhibit two or more of these behaviors. For example, deliberative learning agents improve their reasoning or planning capabilities over time, and reactive service agents perform services in response to environmental changes. A multi-agent system can enable more complex behaviors than monolithic programs (Panait and Luke, 2005), which can lead to powerful emergent behavior (Cucker and Smale, 2007).
An agentic workflow can be formalized as a graph of actions—rather than tasks, as in typical DAG-based workflows—where agents request and perform actions on behalf of one another, enabling dynamic coordination and the collective pursuit of complex, distributed objectives. Let the environment $E$ represent the external state space, influencing and influenced by the actions of agents and other entities. Agents in the environment are represented by a deployment $d(A,R):A\rightarrow R$ of agents $a\in A$ on to resources $r\in R$. Each agent implements a behavior, and an agent that knows the behavior of a peer agent can request the peer to perform an action through message passing. Thus, there exists a directed graph representing the peer relationships between the agents; an edge $e=(a_{i},a_{j})$ in this graph implies that $a_{i}$ knows of $a_{j}$ and can request actions from $a_{j}$. Sink nodes in the graph represent agents that only perform atomic actions, whereas source nodes may represent deliberative or reactive agents that trigger actions on other agents. A cut vertex (articulation point) in the graph can represent an agent that serves as an interface or gateway to another, possibly more complex, multi-agent system.
The deployment of agents can execute workflows. An agentic workflow $W$ is modeled as a directed graph where nodes are a tuple $(x,a)$ of an action to perform and the agent performing the action, and edges representing the source agent and action that triggered the subsequent action. A workflow is typically implicitly encoded within the agent behaviors of a multi-agent system. I.e., the graph $W$ is not explicitly materialized and agents do not need to know $W$ in order to execute. An agent only needs to be concerned with its local view of executing requested actions and requesting actions from peers. Thus, workflows may often be highly dynamic as agent behaviors react to changing states.
3. Academy Design
Designing a middleware that can express the diverse demands of scientific applications and leverage federated research infrastructure is challenging. In the design of Academy, we aim to address the following high-level challenges: How to represent, in code, the declaration of and interaction between agents? How to deploy agents across federated infrastructure? How to achieve performance across heterogeneous systems, networks, and storage? Thus, we begin by outlining key requirements, before we introduce the high-level architecture and detailed implementation choices. The name Academy alludes to societies of artists and scholars that, while independent, collaborate and share similar goals.
3.1. Academy Requirements
Writing scientific applications as agentic workflows, rather than using traditional workflow models, can require a considerable shift in conceptual thinking. To reduce developer friction, our design emphasizes familiarity—using well-known programming patterns—and simplicity—providing a small set of primitives and inviting users to invent new patterns and techniques. With these principles in mind, we define requirements in four areas:
  • Representation: Agent behavior must be expressed in code, supporting control loops, actions that can be performed, and local state. Multiple agents may be instantiated with the same behavior. Agents should not share state.
  • Interaction: Agents and clients must be able to interact. They must be able to address a specific agent, perform one or more actions, modify local state, or create and terminate agents.
  • Communication: Agents and clients communicate asynchronously and are temporally decoupled (e.g., a message sent to an offline agent should be read when that agent is next online). Agents may be deployed in diverse environments with heterogeneous network environments (e.g., asymmetric networks and firewalls restricting connections).
  • Execution: An agent performs actions in response to requests from clients or other agents. Agent control loops may run in perpetuity or exit before the end of the agent’s lifetime. Agents may be launched using different mechanisms that are dependent on the application or environment.
These requirements are an extension of actor systems; therefore, our system inherits properties including simplified concurrency, message processing ordering, loose coupling, error isolation, and modularity (Hewitt et al., 1973).
Our implementation focuses on mechanism rather than policy. That is, we discuss how applications can use Academy to achieve certain outcomes without prescribing what agents, or more generally, applications should do. This avoids constraining, or worse, alienating possible use cases and results in a flexible framework suitable for solving many disparate problems. Further, we describe the components within the architecture in terms of abstract interfaces (i.e., without mandating implementation details such as message protocols, state formats, or ordering) to enable further experimentation and optimization, but we still aim to provide implementations that are suitable for most use cases (as demonstrated in the evaluation). Features such as fault tolerance and resilience, resource allocation, and authentication and authorization, while important, are not listed as explicit requirements because applications have varying demands that preclude one-size-fits-all solutions.
3.2. Academy Architecture
Academy is a middleware for expressing agentic workflows and deploying multi-agent systems across federated resources. Its architecture strongly decouples the implementation of agent behavior from execution and communication to simplify the development of new agents while maintaining flexibility in deployment.
Figure 2. Agents and clients in Academy interact via handles to invoke actions asynchronously. Agents implement a behavior, defined by their actions, control loops, and state. Academy decouples the control and data planes through the launcher and exchange components that manage spawning agents and communication, respectively.
[…]
[Academy system architecture.]
As depicted at a high level in Figure 2, an Academy deployment includes one or more agents and zero or more clients. An agent is a process that executes a behavior, where a behavior is defined by a local state, a set of actions, and a set of control loops. Agents are executed remotely using a launcher. Once running, an agent concurrently executes all of its control loops and listens for messages from clients, which can be other agents or programs.
A client interacts with an agent through a handle, a term we borrow from actor frameworks. A handle acts like a reference to the remote agent and translates method calls into action request messages. Each entity (i.e., client or agent) has an associated mailbox that maintains a queue of messages sent to that entity by other entities. Mailboxes are maintained by an exchange such that any client with access to a given exchange can send messages to the mailbox of another agent in the exchange and receive a response through its own mailbox.
3.3. Academy Implementation Details
Academy is implemented as an open-source Python library, available on GitHub and PyPI. [^1] https://github.com/proxystore/academy We target Python for its broad compatibility with scientific workflow codes and libraries, but both the architecture and individual components could be implemented in other languages.
1import time, threading
2from academy.behavior import Behavior, action, loop
3
4class Example(Behavior):
5 def __init__(self) -> None:
6 self.count = 0 # State stored as attributes
7
8 @action
9 def square(self, value: float) -> float:
10 return value**2
11
12 @loop
13 def count(self, shutdown: threading.Event) -> None:
14 while not shutdown.is_set():
15 self.count += 1
16 time.sleep(1)
Listing 1: Example agent behavior definition.
3.3.1. Behavior
An agent behavior is implemented as a Python class that inherits from the base Behavior type, as shown in 1. This class-based approach is simple, so existing code can be easily transformed into agents, and extensible through inheritance and polymorphism. Instance attributes maintain the agent’s state, and methods define the actions and control loops.
The @action decorator marks a method as an action, allowing other entities to invoke it remotely. (In the future, we plan to support adding metadata to the @action behavior to aid discovery discussed in Section 3.3.4.) A behavior can invoke actions on itself, as actions are simply Python methods. Methods not decorated as @action are private to the behavior. The @loop decorator marks a method as a control loop. Control loops are executed in separate threads, so a shared threading.Event is passed as an argument to each loop that signals when the agent is shutting down so that control loops can gracefully exit. A control loop can terminate early and the agent will remain running. Commonly, control loops are used to execute a routine on a regular interval, such as to check the state of the environment, or in response to an event. We provide two special control loop decorators, @timer and @event, that simplify behavior implementations for these scenarios.
Two special methods, on_setup() and on_shutdown(), allow behaviors to define callbacks when starting or shutting down, such as to load/store state or initialize/destroy resources. Multiple inheritance of behaviors enables the creation of composite agents.
3.3.2. Agent
An Agent is a multithreaded entity that executes a behavior and manages communication with other entities. It is instantiated with a behavior, unique identifier (the address of the agent’s mailbox in the exchange), and exchange interface. An agent is a callable object that when run: (1) invokes the on_setup() callback of the behavior, (2) starts each @loop method in a separate thread, (3) spawns a thread to listen for new messages in the agent’s mailbox, and (4) waits for the agent to be shut down. An @action method is executed in a thread pool when requested remotely so as to not block the handling of other messages. Behaviors can optionally specify the maximum action concurrency.
Agents are designed to be long-running, but can be terminated by sending a shutdown request. Upon shutdown, the shutdown Event, passed to each @loop, is set; running threads are instructed to shutdown and waited on; and the on_shutdown() callback is invoked. Alternatively, an agent can terminate itself by setting the shutdown event. Similarly, an exception raised in an @loop method will shutdown the agent by default but can optionally be suppressed to keep the agent alive. Exceptions raised when executing @action methods are caught and returned to the remote caller.
The use of multi-threading means that behavior implementations must be aware of the caveats of Python’s global interpreter lock (GIL). Compute-heavy actions can dispatch work to other parallel executors, such as process pools, Dask Distributed (Rocklin, 2015), Parsl (Babuji et al., 2019), or Ray (Moritz et al., 2018). We discuss these patterns in more detail in Section 3.4.4. In the future, we would like to support async behaviors and exchanges for improved I/O performance, but scientific computing libraries in Python are not typically async compatible. In Python 3.13 and later, we provide experimental support for free-threading builds, which disable the GIL, enabling full multi-core performance. At this time, however, third-party library support for free-threading builds is limited.
Our decision to decouple behavior definitions from agent execution is deliberate. As behaviors encode application-level logic, we want them to be easily testable and reusable, independent of deployment details. Existing code bases can trivially transition an existing class definition into an agent by inheriting from Behavior and decorating with @action as needed, and behavior classes can still be used independently (i.e., not as a running agent).
3.3.3. Handles
Interacting with an agent is asynchronous; an entity sends a message to the agent’s mailbox and waits to receive a response message in its own mailbox. A handle is a client interface to a remote agent used to invoke actions, ping, and shutdown the agent. Each handle acts as a reference to that agent, translating each method call into a request message that is sent via the exchange and returning a Future. The handle also listens for response messages and accordingly sets the result on the appropriate Future. Rather than creating a return mailbox and listener thread for each handle that a client or agent may have, Academy will multiplex communication for multiple handles within a single process through a single mailbox. This multiplexing ensures that only one mailbox listener thread is needed per process (i.e., agent or client).
3.3.4. Exchange
Entities communicate by sending and receiving messages to and from mailboxes. An exchange hosts these mailboxes, and the Exchange protocol defines the interface to an exchange. Namely, the Exchange defines methods for registering new agent or client mailboxes, sending and receiving messages, and creating handles to remote agents. Registering an agent or client involves creating a unique ID for the entity, which is also the address of its mailbox, and initializing that mailbox within the exchange.
A mailbox has two states: open and closed. Open indicates that the entity is accepting messages, even if, for example, an agent has not yet started or is temporarily offline. Closed indicates permanent termination of the entity and will cause MailboxClosedError to be raised by subsequent send or receive operations to that mailbox.
Exchanges also provide mechanisms for agent discovery by querying based on agent behaviors. This also works with superclasses of behaviors. Consider, for example, a behavior ProteinFolder that can fold proteins (Anfinsen, 1973) and another behavior OpenProteinFolder that inherits from ProteinFolder and specifically uses the OpenFold model (Ahdritz et al., 2024). Querying for ProteinFolder would return the IDs of all agents inheriting from ProteinFolder whereas querying for OpenProteinFolder would return only specific agents using the OpenFold model. In the future, agents could provide additional metadata to enhance discovery.
Users can define custom exchanges to address specific hardware or application characteristics. We provide two exchange implementations for local and distributed agent deployments. The thread exchange stores messages in-memory and is suitable for agents running in separate threads of a single process, such as when testing.
The distributed exchange enables communication between entities across wide-area networks. Core to the distributed exchange is an object store that persists information about registered entities. A hybrid approach is used for message passing: direct messaging is preferred, and indirect message passing via the object store is available as a fallback. Upon startup, an entity writes its location (i.e., address and port) to the object store; peers that want to send a message can attempt to send directly to the entity’s address. If the peer is offline or a direct connection fails, such as in the presence of NAT (network address translation) or firewall restrictions, messages are appended to the list of pending messages in the object store. Entities continuously listen to incoming messages from peers and pending messages in the object store. Entities cache successful communication routes locally to reduce queries to the object store. Our implementation use TCP (transmission control protocol) sockets for direct messaging and a Redis server as the object store. Redis provides low-latency communication and optional replication, but applications that need greater fault-tolerance could consider DHT-based (distributed hash table) object stores.
We optimize the exchange for low latency, as control messages are typically small: $O$(100) bytes. However, action request and response messages can contain arbitrarily sized serialized values for arguments and results that can induce considerable overheads when messages are sent indirectly via the object store. To alleviate these overheads, we pass large values by reference and perform out-of-band data transfers by using ProxyStore (Pauloski et al., 2023, 2025), which provides pass-by-reference semantics in distributed computing through proxy objects. Proxy objects act like references (cheap to serialize and communicate) and automatically de-reference themselves to the true object using performant data storage and communication methods. For example, ProxyStore can leverage RDMA (remote direct memory access) transfers via Mochi (Ross et al., 2020) and UCX (Shamis et al., 2015), GridFTP via Globus Transfer (Chard et al., 2014), and reliable peer-to-peer UDP (user datagram protocol) through NAT hole-punching. ProxyStore also provides two key optimizations useful within Academy: proxies can be forwarded to actions executed on other agents without incurring additional data transfers and proxies can be asynchronously resolved to overlap communication and computation.
3.3.5. Launcher
An agent can be run manually, but the intended method of execution is via a launcher, which manages the initialization and execution of agents on remote resources. The Launcher protocol defines a launch() method with parameters for the behavior, exchange, and agent ID and returns a handle to the launched agent. Users can create custom implementations; we provide the following four that cover most use cases:
  • Thread: Runs agents in separate threads of the same process. Useful for local development and testing or for light-weight or I/O bound agents.
  • Process: Runs agents in separate processes on same machine.
  • Parsl: Runs agents across the workers of a Parsl Executor (Babuji et al., 2019). Parsl supports execution across local, remote, and batch compute systems.
  • Globus Compute: Runs agents across Globus Compute Endpoints (Chard et al., 2020). Globus Compute is a cloud-managed function-as-a-service (FaaS) platform which can execute Python functions across federated compute systems.
The last three launchers support mechanisms to automatically restart agents if they exit unexpectedly. It is common for different agents in an application to be executed with different launchers, but all agents must be registered to the same exchange to interact.
1from academy.exchange.thread import ThreadExchange
2from academy.launcher.thread import ThreadLauncher
3from academy.manager import Manager
4
5with Manager(
6 exchange=ThreadExchange(), # Can be swapped with
7 launcher=ThreadLauncher(), # other implementations
8) as manager:
9 behavior = Example() # From Listing 1
10 handle = manager.launch(behavior)
11
12 future = handle.square(2)
13 assert future.result() == 4
14
15 handle.shutdown() # Or via the manager
16 manager.shutdown(handle.agent_id, blocking=True)
Listing 2: Example of initialization, spawning, using, and shutting down an agent using the Manager interface.
A Manager combines an exchange and one or more launchers to provide a single interface for launching, using, and managing agents. Each manager has a single mailbox in the exchange and multiplexes that mailbox across handles to all of the agents that it manages. This reduces boilerplate code, improves communication efficiency, and ensures stateful resources and threads are appropriately cleaned up. An end-to-end example is provided in 2.
3.4. Common Patterns
We have introduced basic building blocks necessary to build multi-agent systems and deploy agents across federated infrastructure. Now we discuss several common patterns that highlight features of Academy and guide users in building new agentic workflows.
3.4.1. State Checkpoints
Research infrastructure can fail; thus, agents may want to perform periodic state checkpointing. The framework does not enforce a specific checkpointing mechanism, as the format, location, and frequency of checkpoints are highly application specific, but on_startup() callbacks can be used to restore state automatically. For convenience, we provide a State API that provides a dictionary-like interface and persists values to the local file system.
3.4.2. Migration
Research infrastructure is typically static, so Academy does not require that the launcher provide mechanisms for automatic agent migration. Some launchers, such as Parsl, will restart agents on different workers if node-level failures cause agents to crash. Applications can also manually migrate agents across different launchers using agent shutdown and checkpointing mechanisms. These features are sufficient for users to implement custom launchers that enable automatic migration, such as to load-balance across resource pools.
3.4.3. Agent Hierarchies
Agents may dynamically need to create and manage child agents, either to offload tasks or to access new behaviors. A parent agent can create new child agents by using the same launcher used to create the parent, or by creating a new launcher. The use of different launchers is common in scenarios where parent agents want to initialize a local multi-agent system. For example, a client may launch an initial set of agents across federated resources using Globus Compute, and then those initial agents spawn more agents on local resources through Parsl.
3.4.4. Resource Pools
High-performance workflows may need to distribute work across many computers. In an agentic model, resource pool allocation can take two forms: agent managed resource pools or agents as resource pools. In the former, an agent allocates a pool of resources using a parallel computing framework, such as Parsl or Ray, and the agent’s actions dispatch work to resources in the pool. In the second pattern, we deploy identical agents across a set of resources and then route action requests across this agent pool (akin to worker pools in HTTP frameworks).
3.4.5. Process-as-a-Service
FaaS systems, such as Globus Compute, provide optimized execution of short-lived, stateless, and ephemeral functions. Academy agents can extend FaaS systems with process-as-a-service capabilities (Copik et al., 2022), enabling applications to utilize longer-lived, stateful, and isolated processes on-demand.
4. Evaluation
We studied the performance characteristics of Academy to answer key questions including: How well does the system scale? How fast can agents be deployed? What is the messaging latency? We also make comparisons to Dask and Ray, two popular frameworks with support for distributed actors in Python. Although Academy agents provide a superset of features provided by actors, these evaluations contextualize the performance of the framework. In these comparisons, we use the terms agent and actor interchangeably.
We conducted experiments using the Aurora supercomputer at the Argonne Leadership Computing Facility (ALCF), unless otherwise stated. Aurora has 10 6241062410\,62410 624 nodes interconnected by an HPE Slingshot 11 network and a high performance DAOS storage system. Aurora nodes contain two Intel Xeon Max CPUs, each with 52 physical cores and 64 GB of high-bandwidth memory; 512 GB of DDR5 memory per socket; and six 128 GB Intel Data Center Max GPUs. In some cases we also use the Polaris supercomputer at ALCF and the compute-zen-3 nodes of Chameleon Cloud’s CHI@TACC cluster (Keahey et al., 2020). Polaris has $560$ nodes interconnected by an HPE Slingshot 11 network. Polaris nodes contains one AMD EPYC Milan processor with 32 physical cores, 512 GB of DDR4 memory, and four 40 GB NVIDIA A100 GPUs. Each compute-zen-3 node contains two 64-core CPUs and 256 GB memory. Experiments were performed using Python 3.10, AutoGen 0.5.1, Dask 2025.2.0, Globus Compute 3.5.0, Parsl 2025.03.03, and Ray 2.43.0.
4.1. Weak Scaling
We measure weak scaling performance from two aspects: agent startup and action completion time. The object store of the exchange is located on the head node of the Aurora batch job to best match the behavior of Dask and Ray.
Figure 3. (Top) Warm-start time for $n$ agents/actors between Academy (using the Parsl launcher), Dask Actors, and Ray Actors. Ray does not benefit from warm-starts because a new process is spawned for each actor. (Bottom) Time to execute 30 actions per agent/actor (weak scaling). Each action sleeps for 1 s. Note the Academy and Ray lines are overlapped.
[…]
[(Top) Warm-start time for $n$ actors/agents between Academy (using the Parsl launcher), Dask Actors, and Ray Actors. (Bottom) Weak scaling evaluation of action completion time.]
4.1.1. Agent Startup Time
We measure the time to spawn $n$ agents in Figure 3 (top). We pre-warm the worker processes by starting and stopping $n$ agents, then record the average startup time over five runs. Specifically, we measure the time between submitting the first agent to receiving a ping from all agents to ensure that they have finished their startup sequence. We configured Academy to use Parsl’s High-throughput Executor as the launcher. Ray always spawns a new process per actor and thus does not benefit from pre-warmed workers leading to high startup overheads at smaller scales. The cold start time with Academy and Dask is comparable to that of Ray and dominated by loading libraries from the shared file system. With warm starts, Academy starts a single actor in 5.5 ms, 2.8×2.8$\times$2.8 × faster than Dask. Academy scales well, starting $3328$ actors in 7.6 s compared to Dask’s 23.4 s, but Ray demonstrates an advantage at this scale with a 3.2 s startup. Since Academy can leverage many launcher types, applications requiring frequent startup of agents can utilize Parsl for low-latency, and applications launching thousands of long-running agents could use Ray.
4.1.2. Action Completion Time
In Figure 3 (bottom), we execute 30 sleep tasks (1 s) per agent and record the total completion time. We set the maximum concurrency to 1 for all agents to ensure that tasks are processed sequentially. Completion time remains constant for Academy and Ray up to 3328 agents while the performance of Dask degrades starting at 104 actors.
4.2. Academy Distributed Exchange
Next, we study the performance of the distributed exchange.
4.2.1. Data Transfer
Figure 4. (Top) Time for a client to invoke a no-op action on an actor as a function of input and output payload size with different optimizations enabled on the distributed exchange. Two scenarios are considered: client and agent are at the same site (left) and different sites (right). (Bottom) Time for a client to invoke a chain of $n$ actions across $n$ agents with a payload size of 10 MB. Each action in the chain is a no-op that passes the input data along to the next agent, and returns the resulting data. The pass-by-reference optimization reduces communication costs among intermediate actions.
[…]
We first investigate the pass-by-reference and direct communication optimizations of the distributed exchange. In baseline, all message data are communicated indirectly between peers via the exchange’s object store. The object store is located remotely on a Chameleon Cloud node. In pass-by-ref, messages are still communicated with the object store, but action arguments and results are replaced with references using ProxyStore. ProxyStore is configured to use ZeroMQ and ProxyStore’s P2P endpoints for intra-site and inter-site transfer of referenced objects, respectively. In direct, messages are communicated directly between peers, circumventing the cloud-hosted object store; this is only possible when peers are located within the same site.
In Figure 4 (top), we measure the time it takes for a client to invoke a no-op action on an agent as a function of input and output payload size. We compare baseline, pass-by-ref, and direct across two scenarios: Aurora → Aurora, where the client and the agent are located on two different Aurora nodes, and Workstation → Aurora, where the client is located on a personal workstation and the agent is located on an Aurora node. The latencies between the three sites are Aurora to Chameleon: 31 ms; Aurora to Workstation: 12 ms; and Workstation to Chameleon: 42 ms. The workstation is limited to an 800 Mbps internet connection.
We observe that network latency to the exchange object store limits performance at smaller payload sizes ($\leq$ 100 KB). Direct, which is possible only in the intra-site scenario, circumvents these latencies. In both scenarios, pass-by-ref alleviates overheads of data transfer to and from the object store by communicating data directly between the client and agent via ProxyStore. For intra-site transfers, pass-by-ref and direct reduce action latency compared to the baseline by 91.2% and 97.6%, respectively, with 100 MB payloads. For inter-site transfers, pass-by-ref reduces action latency by 78.8%.
Pass-by-ref also reduces overheads when actions pass data to subsequent actions, a common pattern in multi-agent systems. We evaluated this optimization by measuring the round-trip time of action chains in which data are passed through $n$ actions, each invoked on a separate agent, and results are returned through each agent as well. Pass-by-ref reduces the size of messages communicated via the exchange, as indicated by the shallower slope in Figure 4 (bottom). Data are only communicated once to the agent that uses the data (here, the last agent in the chain).
4.2.2. Handle Multiplexing
Figure 5. Maximum no-op action throughput for a single agent requesting actions from $n$ worker agents. The handle multiplexing optimization improves performance by reducing the number of mailbox listener threads from $n$ to 1.
[…]
[Handle multiplexing ablation study.]
As described in Section 3.3.3, the communication of multiple handles within a process is multiplexed through a single mailbox. Without this optimization, each handle held by a client process or agent would create a thread for communication. We evaluated this optimization in Figure 5 by creating one agent that submits a bag-of-tasks to $n$ worker agents and comparing the task throughput with (multiplex) and without (baseline) mailbox multiplexing. Multiplexing improves throughput by 41.7% with 52 worker agents due to reduced threading overheads.
4.3. Agent Messaging
Here, we investigate the performance of agent messaging. As in Section 4.1, the object store of the exchange is located on the head node of the Aurora batch job.
[…]
4.3.1. Action Latency
In Figure 6 (top), we show action latency—the time between sending an action request and receiving a result— between two agents on different nodes. We vary the input/output payload size to understand data transfer overheads. The mean and standard deviation roundtrip latencies are 385$\pm$301 $\mu$s in Academy, 1186$\pm$1059 $\mu$s in Dask, and 526$\pm$308 $\mu$s in Ray for the smallest 10 KB payloads, with latencies increasing with payload size.
4.3.2. Action Throughput
We measure the maximum action throughput for a single agent by submitting a bag of no-op tasks to a pool of worker agents (following the “agents as process pools” pattern from Section 3.4.4). The pool contains 208 agents across four nodes to ensure that each worker agent is not over-saturated with work. That is, the single submitter agent is the limiting factor for performance. Academy, Dask, and Ray achieve maximum throughputs of 3.4K, 185, and 14.1K action/s, respectively. Academy is 18$\times$ faster than Dask but 4$\times$ slower than Ray; however, this is a worst case scenario with no-op tasks and we believe >3K actions/s to be sufficient for real-world agents.
4.3.3. Agent Conversations
In Figure 6 (bottom), we simulate a common pattern in LLM agents where two agents have a back-and-forth conversation. We compare Academy to AutoGen, a popular framework for creating multi-agent AI applications. Each agent is run in a different process on the same node. Agents send ten messages back-and-forth, and we repeat with varying message sizes to simulate different kinds of conversations (i.e., text-only versus multi-modal). AutoGen’s distributed agent runtime uses gRPC which has a maximum message size of 4 MB. Academy has comparatively lower overhead messaging in distributed settings.
4.4. Memory Overhead
Figure 7. Memory used by $n$ agents/actors. We encountered Ray crashes when deploying 104 actors on a single Aurora node (i.e., all cores on both sockets).
[…]
[Memory used by $n$ actors/agents.]
We show memory used as a function of number of agents in Figure 7; for Academy, we compare two launchers: a low-overhead but single-node process-pool and Parsl’s High-throughput Executor. For fairness, we disable features in Dask and Ray that may reduce performance, such as dashboards, and set the initial Ray object store size to the smallest possible value. Academy agents have low memory overheads, making them suitable for memory-constrained devices, such as when deployed across edge devices via the Globus Compute launcher. The Ray cluster head worker has high memory overhead, but that initial overhead is amortized as the number of actors is increased, indicating that each actor has modest overhead.
5. Case Studies
We use three applications to demonstrate the practicality, generality, and robustness of our system in real-world settings. These examples illustrate how Academy integrates with existing research infrastructure, supports a range of computational patterns, and adapts to the varying demands of scientific applications. They validate key design choices, uncover integration challenges, and provide guidance to researchers building agentic workflows.
5.1. Materials Discovery
Figure 8. We execute MOFA by deploying agents across federated infrastructure with Globus Compute. The Assembler, Database, and Estimator run on Chameleon Cloud nodes with fast single-core performance; the Generator and Validator run on Aurora login nodes and execute AI and simulation tasks, respectively, on Aurora compute nodes; and the Optimizer runs on a Polaris login node and executes simulation tasks on Polaris compute nodes. Each agent is responsible for a single MOFA stage, and agents cooperate through message passing, such as to request more work and trigger periodic events. The agents on Aurora and Polaris use Parl to scale resources up and down based on workload needs.
[…]
[Agentic MOFA workflow design.]
MOFA (Yan et al., 2025) is an online learning application for generating, screening, and evaluating metal organic frameworks (MOFs) that couples generative AI methods with computational chemistry. MOFs are polymers composed of inorganic metal clusters and organic ligands; their porosity and large surface area make them suitable for gas adsorption applications such as carbon capture (Furukawa et al., 2013). The goal of MOFA is to generate high-performing candidates by intelligently navigating the intractable combinatorial space of possible MOF structures. MOFA is representative of a broad class of scientific workflows that require careful integration of heterogeneous tasks spanning AI and simulation.
MOFA involves five stages: (1) a generative AI model produces candidate ligands; (2) these ligands are combined with predefined metal clusters to assemble candidate MOFs; (3) the candidates undergo iterative screening and validation using a series of molecular dynamics simulations; (4) CO2 adsorption properties of the most promising structures are simulated and recorded in a database; and (5) the generative model is periodically retrained on the accumulated results to enhance its performance over time. MOFA utilizes Colmena (Ward et al., 2021) to coordinate the flow of data between stages and to distribute computations across CPU and GPU resources within a single batch job. However, this design has key limitations: stages cannot be deployed across heterogeneous resources, such as to leverage hardware best optimized for the specific computations; stages cannot independently scale in or out—resources are bound by the size of a single job; integrating new components within tightly coupled code is challenging; and integration with asynchronous processes, such as synthesis in a real laboratory, are infeasible.
Figure 9. Execution trace of the agentic MOFA workflow of Figure 8 over three hours. (Top) Active tasks per agent. The vertical axis height represents the maximum size of the resource pool allocated by each agent (i.e., CPUs or GPUs). Assembler tasks are short and infrequent. (Middle) Cumulative tasks submitted per agent. (Bottom) Active workers allocated in each agent’s resource pool. Worker allocations vary with demand (as in Assembler and Estimator) or batch job wall times (as in Generator, Validator, and Optimizer).
[…]
[Agentic MOFA workflow execution trace.]
MOFA is an excellent candidate for an agentic workflow, as we demonstrate by porting MOFA to use Academy and deploying the workflow across federated resources: see Figure 8. We express MOFA through six agents: Database, Generator, Assembler, Validator, Optimizer, and Estimator. Each agent is responsible for a different component of the workflow and manages its own resources (i.e., storage and compute). Agents are remotely deployed across Chameleon Cloud nodes and the login nodes of Aurora and Polaris via Globus Compute, and communicate via the distributed exchange backed by a Redis instance in Chameleon Cloud.
An execution trace of the agentic MOFA workflow (Figure 9) shows how each agent scales out its allocated resources as work becomes available, and in the case of Assembler and Estimator, scale down when their workload decreases. The Generator, Validator, and Optimizer consistently have work to do but their batch jobs within which workers run have 60 minute wall times that expire and then must be resubmitted, causing temporary drops the the number of workers. Active tasks that are killed are automatically restarted in the next job. This separation of concerns is key to enabling long-running workflows—resource infrastructure is not persistently available and agents will need to be able to adapt to that varying availability. A second benefit of this model is the loose coupling between agents. For example, the specific implementation of a given agent can be trivially swapped provided the behavior (i.e., the API that agents expose) remains the same. In addition, it becomes easier to integrate future agents, such as to incorporate embodied agents that interact with self-driving labs to synthesize and evaluate the best-performing MOFs in the real-world. While automated MOF synthesis is not yet practical, the capabilities of self-driving labs are rapidly improving (Abolhasani and Kumacheva, 2023; Vescovi et al., 2023), and it is tangible to envision a future where these loosely coupled agentic workflows incorporate services provided by self-driving labs through embodied agents.
5.2. Decentralized Learning
In decentralized machine learning a set of models learn collaboratively across distributed datasets (Hegedűs et al., 2019). This paradigm is particularly relevant today as data are increasingly generated in decentralized settings and transfer to a centralized location can be infeasible for cost and privacy reasons. Each device in a decentralized learning workflow performs three steps: (1) train a model on local data for a set number of iterations; (2) receive models from neighboring devices and send its own model to neighbors; and (3) update the local model via an all reduce operation performed across its own and received models. Reframing the decentralized learning workflow as an agentic workflow is a natural and powerful extension.
We implement a decentralized and asynchronous machine learning exemplar using Academy. The agents and the communication channels between them can be represented as a graph where nodes are agents and edges are communication channels. We choose a powerlaw cluster graph to approximate real-world networks (Holme and Kim, 2002). Each agent is responsible for training its local model, receiving neighboring agents’ models, and aggregating received models with its own model on a periodic basis. Each agent uses a copy of the MNIST dataset (Deng, 2012). The agents are configured to use pass-by-ref with ProxyStore as the transfer backend. Thus data communication between agents follows the network topology. We investigate the cost of distributing updates from all agents as we scale the size of the graph for different model sizes in Figure 10. We do not show training and aggregation time as it is approximately the same for all model sizes and does not increase with the number of agents. The agents are deployed on Aurora using Parsl, where each agent is pinned to a single GPU tile (two tiles per physical GPU), allowing 12 agents per node. Our results demonstrate Academy’s ability to support more than 1500 autonomous agents working collaboratively with no client coordination (as can be seen by the constant time in Figure 10).
Figure 10. Model communication time to an agent’s neighbors averaged over five rounds of decentralized training. Training time and aggregation time are excluded since they are nearly constant.
[…]
5.3. Information Extraction
Figure 11. Execution trace of the agentic MCQ workflow processing 10 manuscripts to generate and validate questions and answers over 15 minutes. The figure shows the active agents and the duration of their tasks. Agents employ either the Mistral-7B-Instruct-v0.3 or Meta-Llama-3-70B-Instruct model, denoted A and B, respectively.
[…]
Exponential growth in scientific publications (Bornmann et al., 2021) creates potential for cross-disciplinary insights that are largely untapped due to the limitations of manual literature review. Automating information extraction from this vast and varied body of work using AI is crucial to accelerate scientific progress. AI methods can be employed to identify and synthesize key findings, methodologies, and datasets across fields and thus to identify connections and facilitate the cross-pollination of ideas that would otherwise go unnoticed (Buchkremer et al., 2019; Sourati and Evans, 2023).
Agentic workflows that leverage LLMs present a transformative new approach to engage with scientific literature. Employing autonomous agents with specific roles and capabilities makes it possible to automate the extraction of information and generation of structured datasets that represent key concepts and findings. Such datasets can be used to fine-tune models and enhance their ability to understand scientific text, answer domain-specific queries, and potentially contribute to tasks like hypothesis generation or literature summarization.
To explore the potential of agentic workflows for thus analyzing the scientific literature we used Academy to implement a system for generating and validating multi-choice questions (MCQs) from research publications (Ch and Saha, 2018; Catlett and Foster, 2025). The workflow includes a PDFParser agent to extract text from a manuscript; two Generator agents that use different LLMs to generate MCQs; an MCQSelector to choose subsets of questions to evaluate; and two MCQAnswerers and two AnswerScorers (again, each with a different LLM) to generate and validate, respectively, answers to questions. The agents use the Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) and Meta-Llama-3-70B-Instruct (at Meta, 2024) models, denoted A and B, respectively.
The beauty of this architecture is that alternative tasks and LLMs are easily integrated by defining new agents; agents can scale up and down in response to demand; and different agents can run concurrently or at different times. We show in Figure 11 an execution trace from a run in which the agents just listed were run concurrently to generate and validate MCQs for 10 publications.
6. Related Work
A workflow is a structured sequence of tasks, typically a directed acyclic graph (DAG), designed to achieve a specific goal, often involving data transformation, analysis, or computational modeling. Frameworks for building workflows take many forms. Parallel computing libraries, such as Dask (Rocklin, 2015) and Ray (Moritz et al., 2018), provide mechanisms for executing functions in parallel across local resources or distributed systems. Similarly, workflow management systems (WMSs) can execute tasks in parallel but also provide mechanisms for defining, optimizing, and monitoring DAG execution (e.g., Airflow (Apache, 2015), Fireworks (Jain et al., 2015), Makeflow (Albrecht et al., 2012), Nextflow (Di Tommaso et al., 2017), Parsl (Babuji et al., 2019), Pegasus (Deelman et al., 2015), Swift (Wilde et al., 2009)). WMSs can be differentiated by how dependency graphs are defined (Pauloski et al., 2024): static configurations files, such as CWL (Crusoe et al., 2022), XML, or YAML; general purpose languages (GPLs); domain specific languages (DSLs); or procedurally through the dynamic execution of a program. The class of workflows supported by these frameworks have two key limitations that we address: tasks are assumed to be pure (i.e., no side-effects) and programs are static, i.e., they cannot adapt to changing environments over time.
Actors are computational entities that enable concurrent computing through message passing (Hewitt et al., 1973). In response to a message, an actor can alter its local state, send messages to other actors, and create new actors. No global state means locks and other synchronization primitives are not required. Actors can enable stateful computations within traditionally stateless programming models, and are supported in parallel computing frameworks (e.g., Akka (Lightbend, 2009), Dask, Orleans (Bernstein et al., 2014), Ray) and function-as-a-service (FaaS) platforms (e.g., Abaco (Garcia et al., 2020), Azure Service Fabric (Microsoft, 2017), PraaS (Copik et al., 2022)). Actor models have been investigated as alternatives for designing computational workflows where communication and coordination are decoupled (Bowers and Ludäscher, 2005). Our system extends the actor model to support autonomous behaviors and federated deployments.
Multi-agent systems can enhance or automate scientific processes. Early work investigated cooperative agent environments for distributed problem solving with minimal human intervention (Drashansky et al., 1995, 1999). Recent work focuses on improving the reasoning capabilities of LLM-backed agents through ontological knowledge graphs and multi-agent systems (Ghafarollahi and Buehler, 2024) and tool-augmented LLMs (Ma et al., 2024). Increasingly popular is the use of multi-agent conversations, in which multiple role-specialized agents interact to collaborate, coordinate, or compete towards goals (Wu et al., 2023). These systems enhance LLM-based tools through better reasoning (Du et al., 2023), validation (Wu et al., 2024), and divergent thinking (Liang et al., 2024), prompting rapid development of frameworks such as LangGraph (LangChain, 2024), Microsoft AutoGen (Wu et al., 2023), OpenAI Swarm (OpenAI, 2024), and Pydantic Agents (Pydantic, 2024). Subsequently, interest in standardizing agent protocols has developed. Anthropic’s Model Context Protocol (MCP) (Anthropic, 2024) defines structured interaction between humans/tools and AI models. Google’s Agent2Agent (A2A) Protocol (Google, 2025) focuses on structured interaction between autonomous agents; each agent serves an HTTP endpoint which is impracticable for many scientific workflows. Multi-agent conversations can proxy scientists in iterative scientific processes—brainstorming ideas, planning experiments, and reasoning about results (Boiko et al., 2023; Wang et al., 2023; Gao et al., 2024; Google Research, 2025)—but these aforementioned systems are designed for local or cloud-native applications and lack the features necessary to deploy agents across federated research infrastructure. We focus on the systems-level challenges of representing and deploying diverse agent types and agentic workflows across heterogeneous environments rather than the applied use of LLMs for workflow steering.
7. Conclusion & Future Work
Advancements in AI, coupled with concurrent advancements in self-driving laboratories, high performance computing, and research data management, open the door for truly autonomous scientific discovery. Realizing this grand vision requires mechanisms for the seamless and dynamic integration of research software and infrastructure. To that end, we introduced Academy, a middleware for developing agentic workflows that engage multi-agent systems spanning federated research infrastructure. This framework enables scalable and flexible orchestration of intelligent agents across heterogeneous resources. We presented solutions to three key challenges: representing and programming agents; communicating among agents; and executing agents across diverse resources. Our evaluations demonstrate that Academy can support high-performance workflows, and three case studies highlight the advantages of agentic workflow design.
In future work, we will explore scoped authentication to control which agents can invoke others, enabling the creation of agent marketplaces where access can be granted, revoked, or delegated. We also plan to expand agent discovery with additional metadata to support AI-steered workflows in which LLMs autonomously identify and use available agents. Recording the relative ordering of agent events (i.e., messages received and state transitions), as in Instant Replay (Leblanc and Mellor-Crummey, 1987), can support provenance within agentic workflows. Last but not least, we will work across scientific research communities to assemble agents for different purposes, and with research facilities to identify obstacles to agent use that may motivate further developments in Academy.
8. Acknowledgments
This research was supported in part by the National Science Foundation under Grants 2004894 and 2209919. We used resources provided by the Argonne Leadership Computing Facility (ALCF), a U.S. Department of Energy (DOE) Office of Science user facility at Argonne National Laboratory supported under Contract DE-AC02-06CH11357, and the Chameleon testbed supported by the National Science Foundation.

### v3

@@
 Empowering Scientific Workflows with Federated Agents
-Abstract.
+Abstract
 Agentic systems, in which diverse agents cooperate to tackle challenging problems, are exploding in popularity in the AI community.
-However, the agentic frameworks used to build these systems have not previously enabled use with research cyberinfrastructure.
+However, existing agentic frameworks take a relatively narrow view of agents, apply a centralized model, and target conversational, cloud-native applications (e.g., LLM-based AI chatbots).
+In contrast, scientific applications require myriad agents be deployed and managed across diverse cyberinfrastructure.
 Here we introduce Academy, a modular and extensible middleware designed to deploy autonomous agents across the federated research ecosystem, including HPC systems, experimental facilities, and data repositories.
@@
 We present microbenchmark results that demonstrate high performance and scalability in HPC environments.
-To demonstrate the breadth of applications that can be supported by agentic workflow designs, we also present case studies in materials discovery, decentralized learning, and information extraction in which agents are deployed across diverse HPC systems.
-Computational Workflows, Distributed Computing, Federated Computing, Multi-Agent Systems, Open-Source Software
-1. Introduction
-The desire to automate scientific processes has led to advancements in many fields, from artificial intelligence (AI) (Zvyagin et al., 2023) and computational workflows (Deelman et al., 2009) to research data management (Allen et al., 2012) and self-driving laboratories (SDL) (Abolhasani and Kumacheva, 2023), but humans typically remain responsible for core aspects of the iterative research cycle, including hypothesis generation, experimental design, code development, and data analysis.
-Often, the human-in-the-loop is the rate-limiting step in discovery.
-This friction increases as the scale and ambition of computational science endeavors grow and leads to inefficient use of research cyberinfrastructure—the federated ecosystem of experimental and observational facilities, data repositories, and high-performance computing (HPC) systems (Miller et al., 2023).
-Figure 1. Cooperative agents, spanning federated research infrastructure (experimental facilities, computational systems, data storage), can enable agentic workflows that autonomously steer discovery.
-[…]
-[Agents deployed across research infrastructure, including experimental facilities, computational systems, and data storage, can facilitate agentic workflows that autonomously engage research infrastructure.]
-Intelligent agents, either as an individual system or composing larger multi-agent systems (MAS), rather than humans, can be the driving entities of discovery.
+To explore the breadth of applications that can be supported by agentic workflow designs, we also present case studies in materials discovery, astronomy, decentralized learning, and information extraction in which agents are deployed across diverse HPC systems.
+11footnotetext: Authors contributed equally to this work.
+I Introduction
+The desire to automate scientific processes has led to advancements in many fields, from artificial intelligence (AI) [1] and computational workflows [2] to research data management [3] and self-driving laboratories (SDL) [4], but humans typically remain responsible for core aspects of the iterative research cycle, including hypothesis generation, experimental design, code development, and data analysis.
+Automated components stall as humans must trigger experiments, manage workflows, correct errors, and make menial decisions.
+This friction increases as the scale and ambition of computational science endeavors grow and leads to inefficient use of experimental and observational facilities, data repositories, and high-performance computing (HPC) systems.
+Intelligent agents, either individually or composing larger multi-agent systems (MAS), rather than humans, can be the driving entities of discovery [5].
 Agents are independent, persistent, stateful, and cooperative—working together to achieve a predefined goal with only intermittent human oversight.
-The contemporaneous explosion of interest in multi-agent systems is largely a consequence of advancements in reasoning capabilities of the large language models (LLMs) often used to back AI agents.
-Expressing components of scientific applications as agents—programs that can perform tasks independently or semi-autonomously on behalf of a client or another agent—is powerful.
+The contemporaneous explosion of interest in multi-agent systems is largely a consequence of advancements in reasoning capabilities of the large language models (LLMs) often used to back AI agents [6, 7, 8].
+Expressing components of scientific applications as agents—programs that can perform tasks independently or semi-autonomously on behalf of a user or another agent—is powerful.
 An agent manages its own local state and exposes a well-defined behavior.
-Agents can perform human roles in iterative scientific processes (Wang et al., 2023) or encapsulate research cyberinfrastructure (e.g., computational resources and procedures, experimental instruments, data repositories) (Gao et al., 2024).
-Significant progress has been made towards developing AI agents that can act on behalf of humans for such tasks as literature synthesis (Lála et al., 2023), hypothesis generation (Google Research, 2025), and data analysis and publication (Ifargan et al., 2025).
-However, existing agent frameworks (e.g., AutoGen (Wu et al., 2023)) are not ready to build and deploy agents that employ federated research cyberinfrastructure.
-New middleware is needed to enable agentic workflows that seamlessly integrate experiment, observation, theory, simulation, AI, analysis, and more, as in Figure 1.
-Frameworks for building agentic workflows are limited in scope and generally target conversational, cloud-native applications (e.g., LLM-based AI chatbots) (LangChain, 2024; Wu et al., 2023; OpenAI, 2024).
-The federated nature of research infrastructure poses unique challenges: distributed resources have diverse access protocols, interactions between computational and experimental entities are asynchronous, and the dynamic availability of resources requires fault-tolerant and adaptive systems.
-Existing frameworks fail to address these intricacies.
-They lack abstractions and mechanisms tailored to support autonomous multi-agent workflows that integrate computation, data management, and experimental control, which leads to brittle, ad hoc integrations that are ill-suited for the demands of modern science.
-Moreover, the inherent complexity of such workflows is compounded by the need to balance efficiency with scientific rigor, especially in applications involving real-time decision-making, iterative exploration, and multi-agent coordination.
-These challenges are often orthogonal and span many levels of abstraction, but achieving this vision where intelligent agents serve as driving entities in scientific discovery requires a paradigm shift in how workflows are designed, orchestrated, and executed.
-We introduce a novel framework for building agentic workflows, emphasizing modularity, statefulness, and interoperability across the diverse research infrastructure.
+Agents can perform human roles in iterative scientific processes [9] or encapsulate research cyberinfrastructure (e.g., computational resources and procedures, experimental instruments, data repositories) [10].
+Figure 1: Cooperative agents, spanning federated research infrastructure (experimental facilities, computational systems, data storage), can enable agentic workflows that autonomously steer discovery.
+However, current agent frameworks (e.g., AutoGen [11]) are not ready to build and deploy agents equipped for scientific applications.
+They are limited in scope and typically apply a centralized model and target conversational, cloud-native applications (e.g., LLM-based AI chatbots) [12, 11, 13].
+Scientific applications, in contrast, typically span geographically (and administratively) separated instruments, supercomputers, robotic facilities, and other resources that collectively constitute federated research infrastructure [14].
+Federation creates unique challenges: distributed resources have diverse access protocols, interactions between resources are asynchronous, and varying resource availability requires fault-tolerant and adaptive systems.
+Existing agentic frameworks are not designed to address these intricacies and workflow management tools designed for federated resources (e.g., Globus Compute [15]) cannot support the long-running and stateful properties of agents natively.
+New middleware is needed to enable agentic workflows that seamlessly integrate experiment, observation, simulation, analysis, and more, as in Figure 1.
+We introduce Academy, a novel actor-based framework for building agentic, scientific workflows, emphasizing modularity, statefulness, and interoperability across federated computing infrastructure.
 Specifically, this work contributes:
   • Academy, a novel, modular, and extensible middleware for expressing agentic workflows and deploying multi-agent systems across federated resources.
-Academy addresses unique challenges in scientific applications, such as high data volumes, variable resource availability, and the heterogeneous nature of experimental and computational systems (Section 3).
-  • Performance analysis of Academy in diverse scenarios yielding insights into the scalability and practical considerations of deploying agentic workflows (Section 4).
-  • Case studies demonstrating the utility of agentic workflow design and highlighting improvements in automation, resource utilization, and discovery acceleration (Section 5).
+Academy addresses unique challenges in scientific applications, such as high data volumes, variable resource availability, and the heterogeneous nature of experimental and computational systems (Section III).
+  • Performance analysis of Academy in diverse scenarios yielding insights into the scalability and practical considerations of deploying agentic workflows (Section IV).
+  • Case studies demonstrating the utility of agentic workflow design and highlighting improvements in automation, resource utilization, and discovery acceleration (Section V).
 These contributions advance the state of the art in multi-agent systems for scientific discovery and establish a foundation for future innovations in autonomous research workflows.
-2. Background
-Agents encompass a rapidly expanding front for AI research, yet agent paradigms can address a breadth of challenges across the computational sciences.
+II Background
+Agents encompass a rapidly expanding front for AI research that can address a breadth of challenges across the computational sciences.
 We begin with a definition of an agent—inspired by prior work—that is sufficiently generic to encompass the various semantic uses of the term.
-Then, we enumerate common high-level classes of agents and formalize agentic workflows, both of which we aim to support in the design of Academy in Section 3.
-An agent is a program that can perform actions independently or semi-autonomously on behalf of a client or another agent.
-This definition is imprecise but presents a powerful conceptual model for distributed computing.
-The agent concept originates from the actor model, a concurrent computing paradigm in which actors encapsulate a local state and communicate through message passing (Hewitt et al., 1973).
+Then, we describe applications that inform the design of Academy, and distill requirements to support agentic scientific workflows.
+An agent is a program that can perform actions independently or semi-autonomously on behalf of a user or another agent.
+This definition incorporates both the modern notion of intelligent LLM agents [11] and the more traditional definitions of agents [16, 17, 18, 19].
+While imprecise, it presents a powerful conceptual model for distributed computing.
+The agent concept originates from the actor model, a concurrent computing paradigm in which actors encapsulate a local state and communicate through message passing [20].
 Agents extend the actor model with the notion of agency—the ability of the agent to engage independently with its environment.
-An agent $a$ is defined by its behavior $B$ and local state $S$.
-The behavior of an agent encompasses a set of actions $x\in X$ (procedures that the agent can perform), and a set of control loops $c\in C$ that define the autonomous behavior (Goodwin, 1995).
-Agents are often long-running, but may also be ephemeral—created to complete a specific task and then exiting.
-Clients and agents can request another agent to perform an action on their behalf through message passing.
-An action can be atomic or composite, invoking other actions on the same or remote agents.
-An agent with actions but no control loops (i.e., $|X|>0$ and $|C|=0$) reduces to an actor.
-Agents come in many flavors.
-Intelligent (deliberative) agents are goal-oriented and reason about what actions to take using internal models and external perception (Wooldridge and Jennings, 1995).
-AI agents, a subset of intelligent agents, use AI methods to make decisions or perform actions.
-In contrast, reactive (observer) agents simply perceive their external environment and react to changes (Nwana, 1996).
-Service agents provide predefined services in response to action requests and come in many forms: resource agents manage and grant access to resources, such as compute or storage, and embodied agents can act in the world, such as through physical actions when paired with a robot body.
-Learning agents adapt their behavior over time to improve performance, often through reinforcement learning (Mnih et al., 2015).
-Composite agents exhibit two or more of these behaviors.
-For example, deliberative learning agents improve their reasoning or planning capabilities over time, and reactive service agents perform services in response to environmental changes.
-A multi-agent system can enable more complex behaviors than monolithic programs (Panait and Luke, 2005), which can lead to powerful emergent behavior (Cucker and Smale, 2007).
-An agentic workflow can be formalized as a graph of actions—rather than tasks, as in typical DAG-based workflows—where agents request and perform actions on behalf of one another, enabling dynamic coordination and the collective pursuit of complex, distributed objectives.
-Let the environment $E$ represent the external state space, influencing and influenced by the actions of agents and other entities.
-Agents in the environment are represented by a deployment $d(A,R):A\rightarrow R$ of agents $a\in A$ on to resources $r\in R$.
-Each agent implements a behavior, and an agent that knows the behavior of a peer agent can request the peer to perform an action through message passing.
-Thus, there exists a directed graph representing the peer relationships between the agents; an edge $e=(a_{i},a_{j})$ in this graph implies that $a_{i}$ knows of $a_{j}$ and can request actions from $a_{j}$.
-Sink nodes in the graph represent agents that only perform atomic actions, whereas source nodes may represent deliberative or reactive agents that trigger actions on other agents.
-A cut vertex (articulation point) in the graph can represent an agent that serves as an interface or gateway to another, possibly more complex, multi-agent system.
-The deployment of agents can execute workflows.
-An agentic workflow $W$ is modeled as a directed graph where nodes are a tuple $(x,a)$ of an action to perform and the agent performing the action, and edges representing the source agent and action that triggered the subsequent action.
-A workflow is typically implicitly encoded within the agent behaviors of a multi-agent system.
-I.e., the graph $W$ is not explicitly materialized and agents do not need to know $W$ in order to execute.
-An agent only needs to be concerned with its local view of executing requested actions and requesting actions from peers.
-Thus, workflows may often be highly dynamic as agent behaviors react to changing states.
-3. Academy Design
-Designing a middleware that can express the diverse demands of scientific applications and leverage federated research infrastructure is challenging.
+II-A Use Cases
+As scientists leverage more advanced computational and experimental resources, and pair them with increasingly capable machine learning models, such as LLMs, the nature of computational science workflows is changing.
+We highlight applications across four emerging patterns that are guiding our development of agentic middleware.
+Steering Applications
+Scientists increasingly want to build applications that delegate the direction and composition of scientific campaigns to ML models.
+Agents simplify the construction of these applications by observing and learning from the results of previous actions and adapting to meet the science goals.
+For example, Colmena is an active learning library for steering simulation campaigns to discover molecules with desirable properties [21].
+In micro-genomics, scientists are using agents to learn the performance of viral identification tools and compose workflows based on sample characteristics [22].
+In physical simulations, agentic systems enable on-the-fly learning that dynamically switches between expensive subroutines and learned surrogate models based on the model uncertainty [23].
+Decomposing Applications across Facilities
+Multi-site workflows are becoming necessary to support complex and heterogeneous scientific applications.
+These applications need to operate autonomously and manage local state while coordinating workflows across administrative domains.
+MOFA [24] is a materials science application designed to discover novel metal organic frameworks (MOFs) for carbon capture.
+It uses a chemical foundation model to generate candidate MOFs and several simulation tools to filter the candidates.
+The tools are best distributed across multiple computing facilities to meet the heterogeneous resource requirements.
+Agents can assist at managing the resources allocated at each facility.
+The Coalition for Epidemic Preparedness Innovations aims to shorten the timeline between disease outbreak and vaccine development to 100 days, which requires coordinating analytics across regions, continuously monitoring outbreaks and trial results, and integrating development and manufacturing into workflows [25].
+Agents can facilitate data, information and resource sharing, and automate components of the workflow.
+Integrating Instruments into Workflows
+As experiments become more data-intensive, scientific applications integrate instruments into computational workflows, or analogously use ML models to control experimental facilities [26].
+Embodied agents can integrate instruments into workflows and provide a mechanism to distribute control to the instrument or experimental site.
+Integral field spectroscopy relies on precise data calibration to study distant spatially complex objects.
+By deploying agents with the instruments, scientists can track the optical parameters necessary for data-processing and accelerate discovery and steering of the telescope [27].
+Self-Driving Laboratories (SDLs) provide an additional tool in chemistry and biology to automatically synthesize materials and measure properties [28].
+Distributing agents to SDLs allows on-site control of experiments and integration of these experiments into computational workflows.
+Conversational Research Assistants
+Scientists are building conversational assistants that help navigate literature, code, documentation, and other data.
+These systems consist of multiple LLM-based agents that interact (e.g., via dialogue or shared context), with different agents being assigned different roles or capabilities.
+For instance, Dr.
+MACS is an astronomy research assistant augmented with retrieval of astronomy literature [29], and ChemGraph is a general purpose chemistry assistant that delegates the construction of chemistry simulation workflows to an LLM [30].
+While constructed today with existing agent frameworks, these applications require access to scientific infrastructure to achieve the autonomy and performance required to meet scientific goals.
+II-B Requirements to Support Federated Agents
+The case studies illuminate benefits of building agentic workflows in science, but also introduce new requirements unmet by existing agent, actor, or workflow systems.
+  1. (R1)
+    Federated Orchestration. Many of the use-cases require launching and managing agents across different computing systems or scientific instruments.
+  2. (R2)
+    Configurable Data Plane.
+Agents may need to frequently communicate at the pace of simulation time-steps on high-performance interconnects or large volumes between facilities and instruments for long running campaigns.
+The communication mechanism that agents use needs to be configurable to adapt to the specific application.
+  3. (R3)
+    Temporally Decoupled Messaging.
+Research infrastructure has varying availability, typically with much lower uptimes than cloud infrastructure [31].
+Communication between agents must cope with facilities being temporarily unavailable.
+  4. (R4)
+    Agent Authentication and Permissions.
+Agents with the capability to use research infrastructure risk exposing the infrastructure to untrusted users.
+Researchers must be able to securely delegate (and revoke) the permissions to use tools and infrastructure to agents.
+  5. (R5)
+    Resilient State Management.
+State is a key feature across use-cases allowing these agentic systems to learn and adapt.
+While state is a necessary feature of the actor model, it is ill supported across workflow management tools that are popular among scientific applications.
+The adoption of specific frameworks is often also influenced by usability requirements such as intuitive and simple representations, local testing and debugging capabilities, etc.
+III AcademyDesign
 In the design of Academy, we aim to address the following high-level challenges: How to represent, in code, the declaration of and interaction between agents?
 How to deploy agents across federated infrastructure?
-How to achieve performance across heterogeneous systems, networks, and storage?
-Thus, we begin by outlining key requirements, before we introduce the high-level architecture and detailed implementation choices.
-The name Academy alludes to societies of artists and scholars that, while independent, collaborate and share similar goals.
-3.1. Academy Requirements
-Writing scientific applications as agentic workflows, rather than using traditional workflow models, can require a considerable shift in conceptual thinking.
-To reduce developer friction, our design emphasizes familiarity—using well-known programming patterns—and simplicity—providing a small set of primitives and inviting users to invent new patterns and techniques.
-With these principles in mind, we define requirements in four areas:
-  • Representation: Agent behavior must be expressed in code, supporting control loops, actions that can be performed, and local state.
-Multiple agents may be instantiated with the same behavior.
-Agents should not share state.
-  • Interaction: Agents and clients must be able to interact. They must be able to address a specific agent, perform one or more actions, modify local state, or create and terminate agents.
-  • Communication: Agents and clients communicate asynchronously and are temporally decoupled (e.g., a message sent to an offline agent should be read when that agent is next online).
-Agents may be deployed in diverse environments with heterogeneous network environments (e.g., asymmetric networks and firewalls restricting connections).
-  • Execution: An agent performs actions in response to requests from clients or other agents.
-Agent control loops may run in perpetuity or exit before the end of the agent’s lifetime.
-Agents may be launched using different mechanisms that are dependent on the application or environment.
-These requirements are an extension of actor systems; therefore, our system inherits properties including simplified concurrency, message processing ordering, loose coupling, error isolation, and modularity (Hewitt et al., 1973).
-Our implementation focuses on mechanism rather than policy.
-That is, we discuss how applications can use Academy to achieve certain outcomes without prescribing what agents, or more generally, applications should do.
-This avoids constraining, or worse, alienating possible use cases and results in a flexible framework suitable for solving many disparate problems.
-Further, we describe the components within the architecture in terms of abstract interfaces (i.e., without mandating implementation details such as message protocols, state formats, or ordering) to enable further experimentation and optimization, but we still aim to provide implementations that are suitable for most use cases (as demonstrated in the evaluation).
-Features such as fault tolerance and resilience, resource allocation, and authentication and authorization, while important, are not listed as explicit requirements because applications have varying demands that preclude one-size-fits-all solutions.
-3.2. Academy Architecture
+How to achieve performance across heterogeneous systems and networks?
+Academy is an open-source Python library, available on GitHub [^1] https://github.com/academy-agents and PyPI.
+We target Python for its broad compatibility with scientific workflow codes and libraries, but both the architecture and individual components could be implemented in other languages.
+III-A Academy Architecture
 Academy is a middleware for expressing agentic workflows and deploying multi-agent systems across federated resources.
 Its architecture strongly decouples the implementation of agent behavior from execution and communication to simplify the development of new agents while maintaining flexibility in deployment.
-Figure 2.
-Agents and clients in Academy interact via handles to invoke actions asynchronously.
+Figure 2: Agents and users in Academy interact via handles to invoke actions asynchronously.
 Agents implement a behavior, defined by their actions, control loops, and state.
-Academy decouples the control and data planes through the launcher and exchange components that manage spawning agents and communication, respectively.
-[…]
-[Academy system architecture.]
-As depicted at a high level in Figure 2, an Academy deployment includes one or more agents and zero or more clients.
-An agent is a process that executes a behavior, where a behavior is defined by a local state, a set of actions, and a set of control loops.
-Agents are executed remotely using a launcher.
-Once running, an agent concurrently executes all of its control loops and listens for messages from clients, which can be other agents or programs.
+Academy decouples the control and data planes through the executor and exchange components that manage spawning agents and communication, respectively.
+As depicted at a high level in Figure 2, an Academy deployment includes one or more agents and zero or more users.
+An agent is a process with a local state, a set of actions, and a set of control loops.
+Agents are executed remotely using an executor.
+Once running, an agent concurrently executes all of its control loops and listens for messages from clients, which can be other agents or users.
 A client interacts with an agent through a handle, a term we borrow from actor frameworks.
 A handle acts like a reference to the remote agent and translates method calls into action request messages.
-Each entity (i.e., client or agent) has an associated mailbox that maintains a queue of messages sent to that entity by other entities.
+Each entity (i.e., user or agent) has an associated mailbox that maintains a queue of messages sent to that entity by other entities.
 Mailboxes are maintained by an exchange such that any client with access to a given exchange can send messages to the mailbox of another agent in the exchange and receive a response through its own mailbox.
-3.3. Academy Implementation Details
-Academy is implemented as an open-source Python library, available on GitHub and PyPI.
-[^1] https://github.com/proxystore/academy We target Python for its broad compatibility with scientific workflow codes and libraries, but both the architecture and individual components could be implemented in other languages.
-1import time, threading
-2from academy.behavior import Behavior, action, loop
+III-B Academy Interaction
+We first describe how scientists program agents and interact with the Academy ecosystem.
+III-B1 Agent Representation
+1from asyncio import Event, sleep
+2from academy.agent import Agent, action, loop
 3
-4class Example(Behavior):
+4class Example(Agent):
 5 def __init__(self) -> None:
@@
 8 @action
-9 def square(self, value: float) -> float:
+9 async def square(self, value: float) -> float:
 10 return value**2
@@
 12 @loop
-13 def count(self, shutdown: threading.Event) -> None:
+13 async def count(self, shutdown: Event) -> None:
 14 while not shutdown.is_set():
 15 self.count += 1
-16 time.sleep(1)
-Listing 1: Example agent behavior definition.
-3.3.1. Behavior
-An agent behavior is implemented as a Python class that inherits from the base Behavior type, as shown in 1.
+16 await sleep(1)
+Listing 1: Example agent definition, showing initialization that sets an internal count variable, an action that squares a supplied value, and a loop that increments the internal count variable once a second.
+An agent is implemented as a Python class that inherits from the base Agent type: see 1.
 This class-based approach is simple, so existing code can be easily transformed into agents, and extensible through inheritance and polymorphism.
@@
 The @action decorator marks a method as an action, allowing other entities to invoke it remotely.
-(In the future, we plan to support adding metadata to the @action behavior to aid discovery discussed in Section 3.3.4.) A behavior can invoke actions on itself, as actions are simply Python methods.
-Methods not decorated as @action are private to the behavior.
-The @loop decorator marks a method as a control loop.
-Control loops are executed in separate threads, so a shared threading.Event is passed as an argument to each loop that signals when the agent is shutting down so that control loops can gracefully exit.
+An agent can invoke actions on itself, as actions are simply Python methods.
+Methods not decorated as @action are private to the agent.
+The @loop decorator marks methods as control loops.
+Control loops extend the actor model to enable the programming of autonomous behavior.
+Control loops are executed concurrently in an event loop; a shared Event is passed to each loop to signal agent shutdown, so that control loops can exit gracefully.
 A control loop can terminate early and the agent will remain running.
 Commonly, control loops are used to execute a routine on a regular interval, such as to check the state of the environment, or in response to an event.
-We provide two special control loop decorators, @timer and @event, that simplify behavior implementations for these scenarios.
-Two special methods, on_setup() and on_shutdown(), allow behaviors to define callbacks when starting or shutting down, such as to load/store state or initialize/destroy resources.
-Multiple inheritance of behaviors enables the creation of composite agents.
-3.3.2. Agent
-An Agent is a multithreaded entity that executes a behavior and manages communication with other entities.
-It is instantiated with a behavior, unique identifier (the address of the agent’s mailbox in the exchange), and exchange interface.
-An agent is a callable object that when run: (1) invokes the on_setup() callback of the behavior, (2) starts each @loop method in a separate thread, (3) spawns a thread to listen for new messages in the agent’s mailbox, and (4) waits for the agent to be shut down.
-An @action method is executed in a thread pool when requested remotely so as to not block the handling of other messages.
-Behaviors can optionally specify the maximum action concurrency.
-Agents are designed to be long-running, but can be terminated by sending a shutdown request.
-Upon shutdown, the shutdown Event, passed to each @loop, is set; running threads are instructed to shutdown and waited on; and the on_shutdown() callback is invoked.
-Alternatively, an agent can terminate itself by setting the shutdown event.
-Similarly, an exception raised in an @loop method will shutdown the agent by default but can optionally be suppressed to keep the agent alive.
-Exceptions raised when executing @action methods are caught and returned to the remote caller.
-The use of multi-threading means that behavior implementations must be aware of the caveats of Python’s global interpreter lock (GIL).
-Compute-heavy actions can dispatch work to other parallel executors, such as process pools, Dask Distributed (Rocklin, 2015), Parsl (Babuji et al., 2019), or Ray (Moritz et al., 2018).
-We discuss these patterns in more detail in Section 3.4.4.
-In the future, we would like to support async behaviors and exchanges for improved I/O performance, but scientific computing libraries in Python are not typically async compatible.
-In Python 3.13 and later, we provide experimental support for free-threading builds, which disable the GIL, enabling full multi-core performance.
-At this time, however, third-party library support for free-threading builds is limited.
-Our decision to decouple behavior definitions from agent execution is deliberate.
-As behaviors encode application-level logic, we want them to be easily testable and reusable, independent of deployment details.
-Existing code bases can trivially transition an existing class definition into an agent by inheriting from Behavior and decorating with @action as needed, and behavior classes can still be used independently (i.e., not as a running agent).
-3.3.3. Handles
-Interacting with an agent is asynchronous; an entity sends a message to the agent’s mailbox and waits to receive a response message in its own mailbox.
+We provide special loop decorators, such as @timer and @event, that simplify agent implementations for common use cases.
+Two special methods, agent_on_startup() and agent_on_shutdown(), allow agents to define callbacks when starting or shutting down, such as to load/store state or initialize/destroy resources.
+Multiple inheritance of agents enables the creation of composite agents.
+III-B2 Agent Invocation
 A handle is a client interface to a remote agent used to invoke actions, ping, and shutdown the agent.
-Each handle acts as a reference to that agent, translating each method call into a request message that is sent via the exchange and returning a Future.
-The handle also listens for response messages and accordingly sets the result on the appropriate Future.
-Rather than creating a return mailbox and listener thread for each handle that a client or agent may have, Academy will multiplex communication for multiple handles within a single process through a single mailbox.
-This multiplexing ensures that only one mailbox listener thread is needed per process (i.e., agent or client).
-3.3.4. Exchange
+Each handle acts as a reference to that agent, translating each method call into a request message that is sent via the exchange and asynchronously waiting on the response message and returning the result.
+The handle decides which mailbox to send from and listen to based on the context where the handle is used; a handle can be passed between users and agents and automatically attaches to the mailbox of the respective client.
+This ensures that there is only one listening task per mailbox, and one mailbox per client (i.e., agent or user).
+III-B3 Manager Class
+A Manager combines an exchange and one or more executors to provide a single interface for launching, using, and managing agents.
+This reduces boilerplate code, improves communication efficiency, and ensures stateful resources and tasks are appropriately cleaned up.
+An end-to-end example is provided in 2.
+III-C Agent Management
+III-C1 Agent Runtime
+Academy provides an agent Runtime that executes an agent and manages communication with other entities.
+It is instantiated with an agent class or instance, a unique identifier (the address of the agent’s mailbox in the exchange), and an exchange interface.
+When started, the Runtime: (1) invokes the agent_on_startup() callback of the agent, (2) spawns tasks for each @loop method, (3) spawns a task to listen for new messages in the agent’s mailbox, and (4) waits for the agent to be shut down.
+We use asyncio to ensure that many @action requests can be handled concurrently without blocking control loop and message listening tasks.
+III-C2 Execution
+An agent can be run manually, but the intended method of deployment is via an executor.
+Any class that implements the concurrent.futures.Executor protocol can be used to launch agents on distributed resources via the Manager interface (R1).
+We use the following executors which cover most local and remote resource types:
+  • Thread: Runs agents in separate threads of the same process. Useful for local development and testing or for lightweight or I/O bound agents.
+  • Process: Runs agents in distinct processes on a machine.
+  • Parsl: Runs agents across the workers of a Parsl Executor [32]. Parsl supports execution on local, remote, and batch compute systems.
+  • Globus Compute: Runs agents across Globus Compute Endpoints [15].
+Globus Compute is a cloud-managed function-as-a-service (FaaS) platform which can execute Python functions across federated compute systems.
+III-C3 State API
+The Academy State API provides a dictionary-like interface for agents to persist in-memory state to storage.
+This API currently supports only writing/reading state to/from disk, but could be extended to other storage modalities.
+Agents can define the agent_on_startup() callback to restore state automatically.
+Given that research infrastructure can fail, agents may want to perform periodic state checkpointing (R5).
+Academy does not enforce a specific checkpointing mechanism, as the format, location, and frequency of checkpoints are highly application specific, agents can use the State API to periodically store checkpoints.
+III-D Agent Communication
 Entities communicate by sending and receiving messages to and from mailboxes.
+The mailboxes serve as mediated communication channels to decouple agents in time (R3); messages persist in a mailbox even when an agent is offline.
 An exchange hosts these mailboxes, and the Exchange protocol defines the interface to an exchange.
-Namely, the Exchange defines methods for registering new agent or client mailboxes, sending and receiving messages, and creating handles to remote agents.
-Registering an agent or client involves creating a unique ID for the entity, which is also the address of its mailbox, and initializing that mailbox within the exchange.
-A mailbox has two states: open and closed.
-Open indicates that the entity is accepting messages, even if, for example, an agent has not yet started or is temporarily offline.
-Closed indicates permanent termination of the entity and will cause MailboxClosedError to be raised by subsequent send or receive operations to that mailbox.
-Exchanges also provide mechanisms for agent discovery by querying based on agent behaviors.
-This also works with superclasses of behaviors.
-Consider, for example, a behavior ProteinFolder that can fold proteins (Anfinsen, 1973) and another behavior OpenProteinFolder that inherits from ProteinFolder and specifically uses the OpenFold model (Ahdritz et al., 2024).
-Querying for ProteinFolder would return the IDs of all agents inheriting from ProteinFolder whereas querying for OpenProteinFolder would return only specific agents using the OpenFold model.
-In the future, agents could provide additional metadata to enhance discovery.
-Users can define custom exchanges to address specific hardware or application characteristics.
-We provide two exchange implementations for local and distributed agent deployments.
-The thread exchange stores messages in-memory and is suitable for agents running in separate threads of a single process, such as when testing.
-The distributed exchange enables communication between entities across wide-area networks.
-Core to the distributed exchange is an object store that persists information about registered entities.
+Namely, the ExchangeClient defines methods for registering new agent or user mailboxes, sending and receiving messages, and creating handles to remote agents.
+Registering an agent or user involves creating a unique ID for the entity, which is also the address of its mailbox, and initializing that mailbox within the exchange.
+A mailbox has two states: active and terminated.
+Active indicates that the entity is accepting messages, even if, for example, an agent has not yet started or is temporarily offline.
+Terminated indicates that the entity is permanently finished and will cause a terminated error to be raised by subsequent send or receive operations to that mailbox.
+Users can define custom exchanges to address specific hardware or application characteristics (R2).
+We provide three exchange implementations: for testing (Thread Exchange), single-site (Hybrid), and distributed (Cloud) deployments.
+III-D1 Thread Exchange
+This implementation stores messages in-memory and is suitable for agents running in separate threads of a single process, such as when testing.
+III-D2 Hybrid Exchange
+This implementation enables communication between entities across local networks (e.g., HPC interconnects).
+It leverages an object store that persists information about registered entities.
 A hybrid approach is used for message passing: direct messaging is preferred, and indirect message passing via the object store is available as a fallback.
 Upon startup, an entity writes its location (i.e., address and port) to the object store; peers that want to send a message can attempt to send directly to the entity’s address.
-If the peer is offline or a direct connection fails, such as in the presence of NAT (network address translation) or firewall restrictions, messages are appended to the list of pending messages in the object store.
+If the peer is offline or a direct connection fails, messages are appended to the list of pending messages in the object store.
+The multi-path design also makes the hybrid exchange suitable for dealing with heterogeneous networks.
 Entities continuously listen to incoming messages from peers and pending messages in the object store.
 Entities cache successful communication routes locally to reduce queries to the object store.
-Our implementation use TCP (transmission control protocol) sockets for direct messaging and a Redis server as the object store.
-Redis provides low-latency communication and optional replication, but applications that need greater fault-tolerance could consider DHT-based (distributed hash table) object stores.
-We optimize the exchange for low latency, as control messages are typically small: $O$(100) bytes.
+Our implementation uses TCP (Transmission Control Protocol) sockets for direct messaging and a Redis server as the object store.
+Redis provides low-latency communication and optional replication, but applications that need greater fault-tolerance could consider (distributed hash table (DHT)-based object stores.
+The security model for the hybrid exchange relies on Redis’s built-in password authentication and access control lists.
+III-D3 Cloud Exchange
+This implementation supports federated agent deployments. It exposes a secure, HTTP-accessible REST API that wraps Redis queues and enforces authorized access to mailboxes (R4).
+The cloud exchange security model is designed to ensure that any communication on the exchange is done by authenticated and permitted agents.
+This guarantees, for instance, that any action request served by an agent comes from an authorized agent/user and similarly response data is only read by the requesting agent/user.
+Authorization and permissions enforcement occur at the exchange server; communication with the exchange is secured using standard TLS.
+Users are authenticated via Globus Auth, an identity and access management platform that supports federated authentication via thousands of supported identity providers [33].
+Each agent is registered as an authenticated entity with the Globus Auth service.
+Creating an agent involves creating a new application identity with Globus Auth; a mailbox on the exchange linked with that identity, and a delegated token for the new agent to authenticate with the exchange; and launching the agent with the delegated token.
+We assume that the launching entity is trusted, and that the launching mechanism is secure.
+When launched, the new agent exchanges a delegated token for a refresh token, allowing long-lived access to the exchange without reauthentication.
+The refresh token is used to obtain access tokens to the exchange; these are introspected by the exchange to verify the client ID matches the expected client of the mailbox, and are cached for 60 seconds.
+Revocation of the delegated token invalidates the refresh token and any derived access token, but the exchange does not respond to the revocation until the cache expires.
+This response time depends on the cache length.
+Users can revoke tokens using the Globus Python SDK or web app.
+Currently, permissions are coarse-grained—an agent is either allowed to communicate or not—but we plan to support finer-grained access controls for agents.
+III-D4 Pass-by-Reference
+We optimize the Hybrid and Cloud exchanges for low latency, as control messages are typically small: OO(100) bytes.
 However, action request and response messages can contain arbitrarily sized serialized values for arguments and results that can induce considerable overheads when messages are sent indirectly via the object store.
-To alleviate these overheads, we pass large values by reference and perform out-of-band data transfers by using ProxyStore (Pauloski et al., 2023, 2025), which provides pass-by-reference semantics in distributed computing through proxy objects.
+To alleviate these overheads, we pass large values by reference and perform out-of-band data transfers by using ProxyStore [34, 35], which provides pass-by-reference semantics in distributed computing through proxy objects.
 Proxy objects act like references (cheap to serialize and communicate) and automatically de-reference themselves to the true object using performant data storage and communication methods.
-For example, ProxyStore can leverage RDMA (remote direct memory access) transfers via Mochi (Ross et al., 2020) and UCX (Shamis et al., 2015), GridFTP via Globus Transfer (Chard et al., 2014), and reliable peer-to-peer UDP (user datagram protocol) through NAT hole-punching.
-ProxyStore also provides two key optimizations useful within Academy: proxies can be forwarded to actions executed on other agents without incurring additional data transfers and proxies can be asynchronously resolved to overlap communication and computation.
-3.3.5. Launcher
-An agent can be run manually, but the intended method of execution is via a launcher, which manages the initialization and execution of agents on remote resources.
-The Launcher protocol defines a launch() method with parameters for the behavior, exchange, and agent ID and returns a handle to the launched agent.
-Users can create custom implementations; we provide the following four that cover most use cases:
-  • Thread: Runs agents in separate threads of the same process. Useful for local development and testing or for light-weight or I/O bound agents.
-  • Process: Runs agents in separate processes on same machine.
-  • Parsl: Runs agents across the workers of a Parsl Executor (Babuji et al., 2019). Parsl supports execution across local, remote, and batch compute systems.
-  • Globus Compute: Runs agents across Globus Compute Endpoints (Chard et al., 2020).
-Globus Compute is a cloud-managed function-as-a-service (FaaS) platform which can execute Python functions across federated compute systems.
-The last three launchers support mechanisms to automatically restart agents if they exit unexpectedly.
-It is common for different agents in an application to be executed with different launchers, but all agents must be registered to the same exchange to interact.
-1from academy.exchange.thread import ThreadExchange
-2from academy.launcher.thread import ThreadLauncher
+For example, ProxyStore can leverage RDMA (remote direct memory access) transfers via Mochi [36] and UCX [37], Globus Transfer [38], and reliable peer-to-peer UDP (user datagram protocol) through NAT hole-punching.
+Two key ProxyStore optimizations are useful within Academy: proxies can be forwarded to actions executed on other agents without incurring additional data transfers and proxies can be asynchronously resolved to overlap communication and computation.
+1from concurrent.futures import ThreadPoolExecutor
+2from academy.exchange.local import LocalExchangeFactory
 3from academy.manager import Manager
 4
-5with Manager(
-6 exchange=ThreadExchange(), # Can be swapped with
-7 launcher=ThreadLauncher(), # other implementations
-8) as manager:
-9 behavior = Example() # From Listing 1
-10 handle = manager.launch(behavior)
-11
-12 future = handle.square(2)
-13 assert future.result() == 4
-14
-15 handle.shutdown() # Or via the manager
-16 manager.shutdown(handle.agent_id, blocking=True)
+5async with await Manager.from_exchange_factory(
+6 # Can be swapped with other implementations
+7 factory=LocalExchangeFactory(),
+8 executors=ThreadPoolExecutor(),
+9) as manager:
+10 agent = Example() # From Listing 1
+11 handle = await manager.launch(agent)
+12
+13 result = await handle.square(2)
+14 assert result == 4
+15
+16 await handle.shutdown() # Or via the manager
+17 await manager.shutdown(handle, blocking=True)
 Listing 2: Example of initialization, spawning, using, and shutting down an agent using the Manager interface.
-A Manager combines an exchange and one or more launchers to provide a single interface for launching, using, and managing agents.
-Each manager has a single mailbox in the exchange and multiplexes that mailbox across handles to all of the agents that it manages.
-This reduces boilerplate code, improves communication efficiency, and ensures stateful resources and threads are appropriately cleaned up.
-An end-to-end example is provided in 2.
-3.4. Common Patterns
-We have introduced basic building blocks necessary to build multi-agent systems and deploy agents across federated infrastructure.
-Now we discuss several common patterns that highlight features of Academy and guide users in building new agentic workflows.
-3.4.1. State Checkpoints
-Research infrastructure can fail; thus, agents may want to perform periodic state checkpointing.
-The framework does not enforce a specific checkpointing mechanism, as the format, location, and frequency of checkpoints are highly application specific, but on_startup() callbacks can be used to restore state automatically.
-For convenience, we provide a State API that provides a dictionary-like interface and persists values to the local file system.
-3.4.2. Migration
-Research infrastructure is typically static, so Academy does not require that the launcher provide mechanisms for automatic agent migration.
-Some launchers, such as Parsl, will restart agents on different workers if node-level failures cause agents to crash.
-Applications can also manually migrate agents across different launchers using agent shutdown and checkpointing mechanisms.
-These features are sufficient for users to implement custom launchers that enable automatic migration, such as to load-balance across resource pools.
-3.4.3. Agent Hierarchies
-Agents may dynamically need to create and manage child agents, either to offload tasks or to access new behaviors.
-A parent agent can create new child agents by using the same launcher used to create the parent, or by creating a new launcher.
-The use of different launchers is common in scenarios where parent agents want to initialize a local multi-agent system.
-For example, a client may launch an initial set of agents across federated resources using Globus Compute, and then those initial agents spawn more agents on local resources through Parsl.
-3.4.4. Resource Pools
-High-performance workflows may need to distribute work across many computers.
-In an agentic model, resource pool allocation can take two forms: agent managed resource pools or agents as resource pools.
-In the former, an agent allocates a pool of resources using a parallel computing framework, such as Parsl or Ray, and the agent’s actions dispatch work to resources in the pool.
-In the second pattern, we deploy identical agents across a set of resources and then route action requests across this agent pool (akin to worker pools in HTTP frameworks).
-3.4.5. Process-as-a-Service
-FaaS systems, such as Globus Compute, provide optimized execution of short-lived, stateless, and ephemeral functions.
-Academy agents can extend FaaS systems with process-as-a-service capabilities (Copik et al., 2022), enabling applications to utilize longer-lived, stateful, and isolated processes on-demand.
-4. Evaluation
-We studied the performance characteristics of Academy to answer key questions including: How well does the system scale?
+IV Evaluation
+We studied the performance characteristics of Academy to answer specific questions: How well does the system scale?
 How fast can agents be deployed?
 What is the messaging latency?
-We also make comparisons to Dask and Ray, two popular frameworks with support for distributed actors in Python.
-Although Academy agents provide a superset of features provided by actors, these evaluations contextualize the performance of the framework.
-In these comparisons, we use the terms agent and actor interchangeably.
+In non-federated settings, we also compare to Dask and Ray, two popular frameworks with support for distributed actors in Python.
 We conducted experiments using the Aurora supercomputer at the Argonne Leadership Computing Facility (ALCF), unless otherwise stated.
-Aurora has 10 6241062410\,62410 624 nodes interconnected by an HPE Slingshot 11 network and a high performance DAOS storage system.
-Aurora nodes contain two Intel Xeon Max CPUs, each with 52 physical cores and 64 GB of high-bandwidth memory; 512 GB of DDR5 memory per socket; and six 128 GB Intel Data Center Max GPUs.
-In some cases we also use the Polaris supercomputer at ALCF and the compute-zen-3 nodes of Chameleon Cloud’s CHI@TACC cluster (Keahey et al., 2020).
+Aurora has 10 62410\,624 nodes interconnected by an HPE Slingshot 11 network and a high performance DAOS storage system.
+Aurora nodes contain two Intel Xeon Max CPUs, each with 52 physical cores and 64 GB of high-bandwidth memory; 512 GB of DDR5 memory per socket; and six 128 GB Intel Data Center Max GPUs (split into two tiles, or logical GPUs each).
+In some cases we also use the Polaris supercomputer at ALCF and the compute-zen-3 nodes of Chameleon Cloud’s CHI@TACC cluster [39].
 Polaris has $560$ nodes interconnected by an HPE Slingshot 11 network.
@@
 Experiments were performed using Python 3.10, AutoGen 0.5.1, Dask 2025.2.0, Globus Compute 3.5.0, Parsl 2025.03.03, and Ray 2.43.0.
-4.1. Weak Scaling
+IV-A Weak Scaling
 We measure weak scaling performance from two aspects: agent startup and action completion time.
-The object store of the exchange is located on the head node of the Aurora batch job to best match the behavior of Dask and Ray.
-Figure 3.
-(Top) Warm-start time for $n$ agents/actors between Academy (using the Parsl launcher), Dask Actors, and Ray Actors.
-Ray does not benefit from warm-starts because a new process is spawned for each actor.
+We use the hybrid exchange with the object store located on the head node of the batch job to best match the behavior of Dask and Ray.
+Figure 3: (Top) Warm-start times for $n$ agents/actors with Academy (with Parsl launcher), Dask Actors, and Ray Actors.
+Ray does not benefit from warm starts because a new process is spawned for each actor.
 (Bottom) Time to execute 30 actions per agent/actor (weak scaling).
@@
 Note the Academy and Ray lines are overlapped.
+IV-A1 Agent Startup Time
 […]
-[(Top) Warm-start time for $n$ actors/agents between Academy (using the Parsl launcher), Dask Actors, and Ray Actors. (Bottom) Weak scaling evaluation of action completion time.]
-4.1.1. Agent Startup Time
-We measure the time to spawn $n$ agents in Figure 3 (top).
-We pre-warm the worker processes by starting and stopping $n$ agents, then record the average startup time over five runs.
-Specifically, we measure the time between submitting the first agent to receiving a ping from all agents to ensure that they have finished their startup sequence.
-We configured Academy to use Parsl’s High-throughput Executor as the launcher.
-Ray always spawns a new process per actor and thus does not benefit from pre-warmed workers leading to high startup overheads at smaller scales.
-The cold start time with Academy and Dask is comparable to that of Ray and dominated by loading libraries from the shared file system.
-With warm starts, Academy starts a single actor in 5.5 ms, 2.8×2.8$\times$2.8 × faster than Dask.
-Academy scales well, starting $3328$ actors in 7.6 s compared to Dask’s 23.4 s, but Ray demonstrates an advantage at this scale with a 3.2 s startup.
-Since Academy can leverage many launcher types, applications requiring frequent startup of agents can utilize Parsl for low-latency, and applications launching thousands of long-running agents could use Ray.
-4.1.2. Action Completion Time
+IV-A2 Action Completion Time
 In Figure 3 (bottom), we execute 30 sleep tasks (1 s) per agent and record the total completion time.
 We set the maximum concurrency to 1 for all agents to ensure that tasks are processed sequentially.
-Completion time remains constant for Academy and Ray up to 3328 agents while the performance of Dask degrades starting at 104 actors.
-4.2. Academy Distributed Exchange
-Next, we study the performance of the distributed exchange.
-4.2.1. Data Transfer
-Figure 4.
-(Top) Time for a client to invoke a no-op action on an actor as a function of input and output payload size with different optimizations enabled on the distributed exchange.
-Two scenarios are considered: client and agent are at the same site (left) and different sites (right).
-(Bottom) Time for a client to invoke a chain of $n$ actions across $n$ agents with a payload size of 10 MB.
-Each action in the chain is a no-op that passes the input data along to the next agent, and returns the resulting data.
-The pass-by-reference optimization reduces communication costs among intermediate actions.
+Completion time remains constant for Academy and Ray up to 3328 agents while Dask performance degrades starting at 104 actors.
+IV-B Academy Exchange
+Figure 4: Time for a client to invoke a no-op action on an actor as a function of input and output payload size with different optimizations enabled on the hybrid and cloud exchanges.
+Two scenarios are considered: client and agent are at the same site using the hybrid exchange (left) and different sites using the cloud exchange (right).
+Next, we investigate performance and optimizations of the hybrid and cloud exchange implementations.
+In baseline, all message data are communicated indirectly between peers via the exchange’s object store.
+For node to node communication on Aurora (Aurora → Aurora), the object store of the hybrid exchange is located on the head node of the Aurora batch job.
+For remote communication, (Workstation → Aurora), the cloud exchange is deployed on AWS.
+In pass-by-ref, messages are still communicated via the object store, but action arguments and results are replaced with references using ProxyStore.
+ProxyStore is configured to use ZeroMQ and P2P endpoints for intra-site and inter-site transfer of referenced objects, respectively.
+In direct, messages are communicated directly between peers, circumventing the object store; this is only possible when peers are located within the same site.
+In Figure 4, we measure the time it takes for a client to invoke a no-op action on an agent as a function of payload size.
+We compare baseline, pass-by-ref, and direct across two scenarios: Aurora → Aurora, where the client and the agent are located on different Aurora nodes and messages are passed via the hybrid exchange, and Workstation → Aurora, where the client is located on a personal workstation, the agent is on an Aurora node, and messages are passed via the cloud exchange.
+The round trip network latencies are: Aurora compute node to Aurora head node: 0.23 ms; Aurora to AWS: 8.8 ms; Aurora to Workstation: 2.2 ms; and Workstation to AWS: 10.4 ms.
+We observe that network latency to the exchange limits performance at smaller payload sizes ($\leq$ 100 KB).
+Direct, which is possible only in the intra-site scenario, circumvents these latencies.
+In both scenarios, pass-by-ref alleviates overheads of data transfer to and from the exchange by communicating data directly between the client and agent via ProxyStore.
+For intra-site transfers, pass-by-ref and direct reduce action latency compared to the baseline by 91.2% and 97.6%, respectively, with 100 MB payloads.
+For inter-site transfers, pass-by-ref reduces action latency by 45.1%.
+IV-C Agent Messaging
+Here, we investigate the performance of agent messaging.
+Figure 5: (Top) No-op action latency between two agents/actors running on separate Aurora nodes versus action input and output payload size.
+(Middle) Action latency between agents located on a workstation and Aurora.
+Academy is compared to Globus Compute, a tool for building federated workflows that only support stateless functions, so state is read from the file-system on every invocation.
+(Bottom) Completion time for a simulated two agent chat where agents send ten messages back and forth with varied message sizes.
+Academy is compared to AutoGen’s distributed runtime.
+IV-C1 Action Latency
 […]
-We first investigate the pass-by-reference and direct communication optimizations of the distributed exchange.
-In baseline, all message data are communicated indirectly between peers via the exchange’s object store.
-The object store is located remotely on a Chameleon Cloud node.
-In pass-by-ref, messages are still communicated with the object store, but action arguments and results are replaced with references using ProxyStore.
-ProxyStore is configured to use ZeroMQ and ProxyStore’s P2P endpoints for intra-site and inter-site transfer of referenced objects, respectively.
-In direct, messages are communicated directly between peers, circumventing the cloud-hosted object store; this is only possible when peers are located within the same site.
-In Figure 4 (top), we measure the time it takes for a client to invoke a no-op action on an agent as a function of input and output payload size.
-We compare baseline, pass-by-ref, and direct across two scenarios: Aurora → Aurora, where the client and the agent are located on two different Aurora nodes, and Workstation → Aurora, where the client is located on a personal workstation and the agent is located on an Aurora node.
-The latencies between the three sites are Aurora to Chameleon: 31 ms; Aurora to Workstation: 12 ms; and Workstation to Chameleon: 42 ms.
-The workstation is limited to an 800 Mbps internet connection.
-We observe that network latency to the exchange object store limits performance at smaller payload sizes ($\leq$ 100 KB).
-Direct, which is possible only in the intra-site scenario, circumvents these latencies.
-In both scenarios, pass-by-ref alleviates overheads of data transfer to and from the object store by communicating data directly between the client and agent via ProxyStore.
-For intra-site transfers, pass-by-ref and direct reduce action latency compared to the baseline by 91.2% and 97.6%, respectively, with 100 MB payloads.
-For inter-site transfers, pass-by-ref reduces action latency by 78.8%.
-Pass-by-ref also reduces overheads when actions pass data to subsequent actions, a common pattern in multi-agent systems.
-We evaluated this optimization by measuring the round-trip time of action chains in which data are passed through $n$ actions, each invoked on a separate agent, and results are returned through each agent as well.
-Pass-by-ref reduces the size of messages communicated via the exchange, as indicated by the shallower slope in Figure 4 (bottom).
-Data are only communicated once to the agent that uses the data (here, the last agent in the chain).
-4.2.2. Handle Multiplexing
-Figure 5.
-Maximum no-op action throughput for a single agent requesting actions from $n$ worker agents.
-The handle multiplexing optimization improves performance by reducing the number of mailbox listener threads from $n$ to 1.
 […]
-[Handle multiplexing ablation study.]
-As described in Section 3.3.3, the communication of multiple handles within a process is multiplexed through a single mailbox.
-Without this optimization, each handle held by a client process or agent would create a thread for communication.
-We evaluated this optimization in Figure 5 by creating one agent that submits a bag-of-tasks to $n$ worker agents and comparing the task throughput with (multiplex) and without (baseline) mailbox multiplexing.
-Multiplexing improves throughput by 41.7% with 52 worker agents due to reduced threading overheads.
-4.3. Agent Messaging
-Here, we investigate the performance of agent messaging. As in Section 4.1, the object store of the exchange is located on the head node of the Aurora batch job.
+IV-C2 Action Throughput
 […]
-4.3.1. Action Latency
-In Figure 6 (top), we show action latency—the time between sending an action request and receiving a result— between two agents on different nodes.
-We vary the input/output payload size to understand data transfer overheads.
-The mean and standard deviation roundtrip latencies are 385$\pm$301 $\mu$s in Academy, 1186$\pm$1059 $\mu$s in Dask, and 526$\pm$308 $\mu$s in Ray for the smallest 10 KB payloads, with latencies increasing with payload size.
-4.3.2. Action Throughput
-We measure the maximum action throughput for a single agent by submitting a bag of no-op tasks to a pool of worker agents (following the “agents as process pools” pattern from Section 3.4.4).
-The pool contains 208 agents across four nodes to ensure that each worker agent is not over-saturated with work.
-That is, the single submitter agent is the limiting factor for performance.
-Academy, Dask, and Ray achieve maximum throughputs of 3.4K, 185, and 14.1K action/s, respectively.
-Academy is 18$\times$ faster than Dask but 4$\times$ slower than Ray; however, this is a worst case scenario with no-op tasks and we believe >3K actions/s to be sufficient for real-world agents.
-4.3.3. Agent Conversations
+IV-C3 Agent Conversations
 In Figure 5 (bottom), we simulate a common pattern in LLM agents where two agents have a back-and-forth conversation.
@@
 Each agent is run in a different process on the same node.
-Agents send ten messages back-and-forth, and we repeat with varying message sizes to simulate different kinds of conversations (i.e., text-only versus multi-modal).
+Agents send 10 messages back and forth, repeating with varying message sizes to simulate different kinds of conversations (i.e., text-only vs.
+multi-modal).
 AutoGen’s distributed agent runtime uses gRPC which has a maximum message size of 4 MB.
 Academy has comparatively lower overhead messaging in distributed settings.
-4.4. Memory Overhead
-Figure 7. Memory used by $n$ agents/actors. We encountered Ray crashes when deploying 104 actors on a single Aurora node (i.e., all cores on both sockets).
-[…]
-[Memory used by $n$ actors/agents.]
-We show memory used as a function of number of agents in Figure 7; for Academy, we compare two launchers: a low-overhead but single-node process-pool and Parsl’s High-throughput Executor.
-For fairness, we disable features in Dask and Ray that may reduce performance, such as dashboards, and set the initial Ray object store size to the smallest possible value.
-Academy agents have low memory overheads, making them suitable for memory-constrained devices, such as when deployed across edge devices via the Globus Compute launcher.
+IV-D Memory Overhead
+Figure 6: Memory used by $n$ agents/actors. We encountered a bug causing Ray to crash when deploying 104 actors on a single Aurora node (i.e., all cores on both sockets).
+We show memory used as a function of number of agents in Figure 6; for Academy, we compare two executors: a low-overhead but single-node process-pool and Parsl’s High-throughput Executor.
+For fairness, we disable features in Dask and Ray that may increase memory, such as dashboards, and set the initial Ray object store size to the smallest possible value.
+Academy agents have low memory overheads, making them suitable for memory-constrained devices, as when deployed to edge devices via Globus Compute.
 The Ray cluster head worker has high memory overhead, but that initial overhead is amortized as the number of actors is increased, indicating that each actor has modest overhead.
-5. Case Studies
-We use three applications to demonstrate the practicality, generality, and robustness of our system in real-world settings.
-These examples illustrate how Academy integrates with existing research infrastructure, supports a range of computational patterns, and adapts to the varying demands of scientific applications.
-They validate key design choices, uncover integration challenges, and provide guidance to researchers building agentic workflows.
-5.1. Materials Discovery
-Figure 8.
-We execute MOFA by deploying agents across federated infrastructure with Globus Compute.
+V Case Studies
+We use four applications to demonstrate the practicality, generality, and robustness of Academy in real-world settings.
+These examples illustrate how Academy integrates with existing research infrastructure, enables distinct capabilities, and adapts to the varying demands of scientific applications.
+More generally, these examples demonstrate how the agent paradigm simplifies the construction of the scientific application patterns identified in Section II.
+V-A Materials Discovery
+Figure 7: MOFA deploys agents across federated infrastructure with Globus Compute.
 The Assembler, Database, and Estimator run on Chameleon Cloud nodes with fast single-core performance; the Generator and Validator run on Aurora login nodes and execute AI and simulation tasks, respectively, on Aurora compute nodes; and the Optimizer runs on a Polaris login node and executes simulation tasks on Polaris compute nodes.
-Each agent is responsible for a single MOFA stage, and agents cooperate through message passing, such as to request more work and trigger periodic events.
-The agents on Aurora and Polaris use Parl to scale resources up and down based on workload needs.
-[…]
-[Agentic MOFA workflow design.]
-MOFA (Yan et al., 2025) is an online learning application for generating, screening, and evaluating metal organic frameworks (MOFs) that couples generative AI methods with computational chemistry.
-MOFs are polymers composed of inorganic metal clusters and organic ligands; their porosity and large surface area make them suitable for gas adsorption applications such as carbon capture (Furukawa et al., 2013).
-The goal of MOFA is to generate high-performing candidates by intelligently navigating the intractable combinatorial space of possible MOF structures.
+Each agent is responsible for a single MOFA stage.
+Agents cooperate by passing messages on the exchange, such as to request more work and trigger periodic events.
+Agents on Aurora and Polaris use Parsl to scale resources up and down based on workload needs.
+MOFA [24] is an online learning application for generating, screening, and evaluating metal organic frameworks (MOFs) that couples generative AI methods with computational chemistry.
+MOFs are polymers composed of inorganic metal clusters and organic ligands that are particularly suitable for gas adsorption applications such as carbon capture [40].
+The goal of MOFA is to generate high-performing candidates by intelligently navigating space of possible MOF structures.
 MOFA is representative of a broad class of scientific workflows that require careful integration of heterogeneous tasks spanning AI and simulation.
 MOFA involves five stages: (1) a generative AI model produces candidate ligands; (2) these ligands are combined with predefined metal clusters to assemble candidate MOFs; (3) the candidates undergo iterative screening and validation using a series of molecular dynamics simulations; (4) CO2 adsorption properties of the most promising structures are simulated and recorded in a database; and (5) the generative model is periodically retrained on the accumulated results to enhance its performance over time.
-MOFA utilizes Colmena (Ward et al., 2021) to coordinate the flow of data between stages and to distribute computations across CPU and GPU resources within a single batch job.
-However, this design has key limitations: stages cannot be deployed across heterogeneous resources, such as to leverage hardware best optimized for the specific computations; stages cannot independently scale in or out—resources are bound by the size of a single job; integrating new components within tightly coupled code is challenging; and integration with asynchronous processes, such as synthesis in a real laboratory, are infeasible.
-Figure 9.
-Execution trace of the agentic MOFA workflow of Figure 8 over three hours.
+TABLE I: Per-node throughputs, in MOFs/min, for the various components of the MOFA application (Figure 7).
+Task Academy Colmena/Parsl
+Machine Throughput Machine Throughput
+Validate Aurora 10.73 Polaris 4.34
+Optimize Polaris 0.85 Polaris 0.85
+Estimate Cloud 6.81 Polaris 1.85
+[…]
+We adapt MOFA to use Academy and deploy the resulting agentic application across federated resources: see Figure 7.
+We use six agents: Database, Generator, Assembler, Validator, Optimizer, and Estimator.
+Each agent is responsible for a different component of the workflow and manages its own resources (i.e., storage and compute).
+Agents are remotely deployed across Chameleon Cloud, Aurora and Polaris.
+Figure 8: Execution trace of the Academy MOFA workflow.
 (Top) Active tasks per agent.
-The vertical axis height represents the maximum size of the resource pool allocated by each agent (i.e., CPUs or GPUs).
-Assembler tasks are short and infrequent.
+The vertical axis height represents the maximum size of the resource pool allocated by each agent.
 (Middle) Cumulative tasks submitted per agent.
@@
 Worker allocations vary with demand (as in Assembler and Estimator) or batch job wall times (as in Generator, Validator, and Optimizer).
-[…]
-[Agentic MOFA workflow execution trace.]
-MOFA is an excellent candidate for an agentic workflow, as we demonstrate by porting MOFA to use Academy and deploying the workflow across federated resources: see Figure 8.
-We express MOFA through six agents: Database, Generator, Assembler, Validator, Optimizer, and Estimator.
-Each agent is responsible for a different component of the workflow and manages its own resources (i.e., storage and compute).
-Agents are remotely deployed across Chameleon Cloud nodes and the login nodes of Aurora and Polaris via Globus Compute, and communicate via the distributed exchange backed by a Redis instance in Chameleon Cloud.
-An execution trace of the agentic MOFA workflow (Figure 9) shows how each agent scales out its allocated resources as work becomes available, and in the case of Assembler and Estimator, scale down when their workload decreases.
-The Generator, Validator, and Optimizer consistently have work to do but their batch jobs within which workers run have 60 minute wall times that expire and then must be resubmitted, causing temporary drops the the number of workers.
-Active tasks that are killed are automatically restarted in the next job.
+An execution trace of the agentic MOFA workflow (Figure 8) shows how each agent scales out its allocated resources as work becomes available, and in the case of the Generator, Validator, and Assembler, scale down when their workload decreases.
+After a ramp up period, the Optimizer consistently has work to do but their batch jobs within which workers run have 60 minute wall times that expire and then must be resubmitted, causing temporary drops in the number of workers.
+Active tasks that are killed are restarted in the next job.
 This separation of concerns is key to enabling long-running workflows—resource infrastructure is not persistently available and agents will need to be able to adapt to that varying availability.
-A second benefit of this model is the loose coupling between agents.
-For example, the specific implementation of a given agent can be trivially swapped provided the behavior (i.e., the API that agents expose) remains the same.
-In addition, it becomes easier to integrate future agents, such as to incorporate embodied agents that interact with self-driving labs to synthesize and evaluate the best-performing MOFs in the real-world.
-While automated MOF synthesis is not yet practical, the capabilities of self-driving labs are rapidly improving (Abolhasani and Kumacheva, 2023; Vescovi et al., 2023), and it is tangible to envision a future where these loosely coupled agentic workflows incorporate services provided by self-driving labs through embodied agents.
-5.2. Decentralized Learning
-In decentralized machine learning a set of models learn collaboratively across distributed datasets (Hegedűs et al., 2019).
-This paradigm is particularly relevant today as data are increasingly generated in decentralized settings and transfer to a centralized location can be infeasible for cost and privacy reasons.
-Each device in a decentralized learning workflow performs three steps: (1) train a model on local data for a set number of iterations; (2) receive models from neighboring devices and send its own model to neighbors; and (3) update the local model via an all reduce operation performed across its own and received models.
-Reframing the decentralized learning workflow as an agentic workflow is a natural and powerful extension.
-We implement a decentralized and asynchronous machine learning exemplar using Academy.
-The agents and the communication channels between them can be represented as a graph where nodes are agents and edges are communication channels.
-We choose a powerlaw cluster graph to approximate real-world networks (Holme and Kim, 2002).
-Each agent is responsible for training its local model, receiving neighboring agents’ models, and aggregating received models with its own model on a periodic basis.
-Each agent uses a copy of the MNIST dataset (Deng, 2012).
+We contrast the Academy implementation of MOFA to a version implemented with Colmena [21], a bespoke simulation campaign framework built on Parsl [32]—a traditional workflow system that does not support stateful federated agents.
+We compare the Academy version, deployed on federated resources, to the Colmena version run as a single batch job on Polaris.
+In addition to using different resources, the Academy execution includes queue wait times to acquire resources while the Colmena execution does not.
+Although the performances are thus not directly comparable, the comparison against Colmena/Parsl (shown in Figure 9) demonstrates how federated agents support more efficient resource utilization.
+With Academy, resource allocations (Figure 8) are managed autonomously by each agent, whereas Colmena allocates resources statically for the entire workflow.
+This, along with the hardware configuration of Polaris, means that when the Colmena workflow performs CPU-only tasks like “estimate,” the GPU resources of the allocation sit idle.
+Figure 9: Resources used for a traditional workflow implementation of MOFA deployed on Polaris.
+The static resource configuration does not match the workload, so in the estimation stage, when tasks are only using the CPU, the rest of the resources are idle.
+A further advantage is that the loose coupling between agents in the Academy model makes it trivial to swap one agent implementation for another, provided that the API exposed by the agent remains the same.
+It also becomes easier to integrate future agents, such as to incorporate embodied agents that interact with self-driving labs to synthesize the best-performing MOFs via physical experiments [4, 28, 41].
+V-B Astronomical Spectroscopy
+Integral field spectroscopy is a powerful technique used in observational astrophysics to gather spatially-resolved spectroscopic data from a target field at once, enabling the study of spatially-complex objects like distant strongly-lensed galaxies—optimal candidates for understanding galaxy formation in the early universe [42].
+Before analyzing the spectra offered by the Integral Field Units for Magellan (IFU-M) instrument, scientists first process and calibrate the data which involves the overarching steps: pre-processing, spatial and wavelength calibrations, and datacube creation.
+A crucial step in pre-processing the data is detecting and filtering cosmic rays through subtracting multiple images of the same patch of sky.
+However, the optics of the telescope shift over time in response to small variations in external factors such as temperature.
+The images must be aligned by finding offset and scaling parameters that minimize the variation between images.
+Previously, the alignment was found through a grid search of parameters between pairs of images.
+We employ a hybrid pattern for this workflow in which Academy conceptually maps the instrument into the workflow by estimating and storing the optical parameters of the telescope (see Figure 10).
+As these parameters vary smoothly in time, the agent uses previously calibrated images to inform the search for new parameters, leading to faster alignment and reduced noise in the processed images.
+The two spectrographs of the instrument are each represented by their own agent.
+Other tasks in the application remain stateless, allowing them to be scheduled to any available worker.
+In the future, deploying the agents on the instrument would allow the agent to use observations (e.g., temperature measurements) to inform the alignment, and provide on-site feedback for targeting observations.
+Figure 10: Execution trace of the IFU-M workflow.
+Each color corresponds to a different task type.
+Tasks within the Parsl workflow are stateless, but invoke actions on agents that track the instrument state.
+V-C Decentralized Learning
+In decentralized machine learning a set of models learn collaboratively across distributed datasets [43].
+This paradigm is particularly relevant today as data are generated in decentralized settings and transfer to a centralized location can be infeasible for cost and privacy reasons.
+Each device in a decentralized learning workflow performs three steps: (1) train model on local data for a set number of iterations; (2) receive models from neighboring devices and send its own model to neighbors; and (3) update the local model via an all-reduce operation performed across its own and received models.
+It is straightforward to reframe such a decentralized learning workflow as an agentic workflow.
+We represent the application as a graph in which nodes are agents and edges are communication channels.
+Each agent trains its local model, receives neighboring agents’ models, and periodically aggregates received models with its own model.
+Using Academy, we simulate decentralized learning on Aurora.
+For the connectivity between devices, we choose a power-law cluster graph to approximate real-world networks [44].
+Each agent uses a copy of the MNIST dataset [45].
 The agents are configured to use pass-by-ref with ProxyStore as the transfer backend.
-Thus data communication between agents follows the network topology.
-We investigate the cost of distributing updates from all agents as we scale the size of the graph for different model sizes in Figure 10.
-We do not show training and aggregation time as it is approximately the same for all model sizes and does not increase with the number of agents.
-The agents are deployed on Aurora using Parsl, where each agent is pinned to a single GPU tile (two tiles per physical GPU), allowing 12 agents per node.
-Our results demonstrate Academy’s ability to support more than 1500 autonomous agents working collaboratively with no client coordination (as can be seen by the constant time in Figure 10).
-Figure 10. Model communication time to an agent’s neighbors averaged over five rounds of decentralized training. Training time and aggregation time are excluded since they are nearly constant.
-[…]
-5.3. Information Extraction
-Figure 11.
-Execution trace of the agentic MCQ workflow processing 10 manuscripts to generate and validate questions and answers over 15 minutes.
+We investigate the cost of distributing updates from all agents as we scale the size of the graph for different model sizes: see Figure 11.
+The agents are deployed on Aurora using Parsl, with each agent pinned to a single GPU tile (two tiles per physical GPU), allowing 12 agents per node.
+Our results demonstrate Academy’s ability to support more than 1500 autonomous agents working collaboratively with no client coordination, as shown by the constant times seen in Figure 11.
+Figure 11: Model communication time to an agent’s neighbors averaged over five rounds of decentralized training. Training time and aggregation time are excluded since they are nearly constant.
+V-D Information Extraction
+Figure 12: Execution trace of the agentic MCQ workflow processing 10 manuscripts to generate and validate questions and answers over 15 minutes.
 The figure shows the active agents and the duration of their tasks.
 Agents employ either the Mistral-7B-Instruct-v0.3 or Meta-Llama-3-70B-Instruct model, denoted A and B, respectively.
-[…]
-Exponential growth in scientific publications (Bornmann et al., 2021) creates potential for cross-disciplinary insights that are largely untapped due to the limitations of manual literature review.
+Exponential growth in scientific publications [46] creates potential for cross-disciplinary insights that are largely untapped due to the limitations of manual literature review.
 Automating information extraction from this vast and varied body of work using AI is crucial to accelerate scientific progress.
-AI methods can be employed to identify and synthesize key findings, methodologies, and datasets across fields and thus to identify connections and facilitate the cross-pollination of ideas that would otherwise go unnoticed (Buchkremer et al., 2019; Sourati and Evans, 2023).
+AI methods can be employed to identify and synthesize key findings, methodologies, and datasets across fields and thus to identify connections and facilitate the cross-pollination of ideas that would otherwise go unnoticed [47, 48].
 Agentic workflows that leverage LLMs present a transformative new approach to engage with scientific literature.
@@
 Such datasets can be used to fine-tune models and enhance their ability to understand scientific text, answer domain-specific queries, and potentially contribute to tasks like hypothesis generation or literature summarization.
-To explore the potential of agentic workflows for thus analyzing the scientific literature we used Academy to implement a system for generating and validating multi-choice questions (MCQs) from research publications (Ch and Saha, 2018; Catlett and Foster, 2025).
+To explore the potential of agentic workflows for thus analyzing the scientific literature, we used Academyto implement a system for generating and validating multi-choice questions (MCQs) from research publications [49, 50].
 The workflow includes a PDFParser agent to extract text from a manuscript; two Generator agents that use different LLMs to generate MCQs; an MCQSelector to choose subsets of questions to evaluate; and two MCQAnswerers and two AnswerScorers (again, each with a different LLM) to generate and validate, respectively, answers to questions.
-The agents use the Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) and Meta-Llama-3-70B-Instruct (at Meta, 2024) models, denoted A and B, respectively.
+The agents use the Mistral-7B-Instruct-v0.3 [51] and Meta-Llama-3-70B-Instruct [52] models, denoted A and B, respectively.
 The beauty of this architecture is that alternative tasks and LLMs are easily integrated by defining new agents; agents can scale up and down in response to demand; and different agents can run concurrently or at different times.
 We show in Figure 12 an execution trace from a run in which the agents just listed were run concurrently to generate and validate MCQs for 10 publications.
-6. Related Work
+VI Discussion
+We describe several lessons learned implementing Academy and applying it to a diverse range of applications.
+Scientific agentic systems require federation and statefulness.
+Science is often collaborative and multi-institutional.
+Research infrastructure (e.g., instruments, HPC systems, etc.) mirrors this inherently federated ecosystem.
+Building intelligent scientific systems requires agents that run in multiple locations and collaborate across institutions.
+Existing workflow systems (e.g., Parsl, Ray, Dask, Globus Compute) either do not support the capabilities necessary for agentic science, or face significant barriers when deployed on research infrastructure.
+Similarly, conversational agentic frameworks (e.g., LangChain, PydanticAI) are not designed for such environments, but can leverage Academy to deliver these capabilities.
+Start-up barriers hinder adoption and deployment.
+Academy initially required that users deploy the hybrid exchange to allow inter-site communication.
+This requirement hindered the transition from local development to multi-site deployment.
+The introduction of the cloud exchange reduced barriers and thus enabled significant improvements (e.g., agent sharing and reuse).
+It also simplified integration with existing execution and data transfer solutions (e.g., Globus Compute, ProxyStore), further reducing barriers of using Academy.
+Agentic middleware should not bias the behaviors that users create.
+Commonly used agentic systems such as LangChain and SmolAgents shoehorn applications into specific patterns (e.g., LLMs with tool-calling).
+Scientific agents behave in diverse ways that often do not fit such patterns.
+Rather than design specific behaviors into the system, we implemented patterns in Academy (e.g., agent hierarchies, agents as resource managers) and reused them in different contexts.
+LLMs combined with tool calling cannot currently construct complex real-world science applications.
+For instance, in the MOFA workflow we experimented with the use of LLMs for such functions as improving MOF design and deciding which simulators and simulations to run.
+However, we found that the large number of tools and tool parameters, often with limited documentation, and sometimes subtle differences among tools, led to frequent problems including improperly configured/invoked codes that ran without errors but produced incorrect results.
+Manual construction of both tools (as Academy actions) and workflow architecture (by invoking handles) informed by domain expertise, created more reliable applications.
+Autonomy and adaptability can improve application performance and resource utilization.
+Static resource allocations often inadequately match the dynamic nature of scientific workloads.
+Agentic systems enable not only component autonomy but also dynamic resource utilization at fine granularity.
+VII Related Work
 A workflow is a structured sequence of tasks, typically a directed acyclic graph (DAG), designed to achieve a specific goal, often involving data transformation, analysis, or computational modeling.
 Frameworks for building workflows take many forms.
-Parallel computing libraries, such as Dask (Rocklin, 2015) and Ray (Moritz et al., 2018), provide mechanisms for executing functions in parallel across local resources or distributed systems.
-Similarly, workflow management systems (WMSs) can execute tasks in parallel but also provide mechanisms for defining, optimizing, and monitoring DAG execution (e.g., Airflow (Apache, 2015), Fireworks (Jain et al., 2015), Makeflow (Albrecht et al., 2012), Nextflow (Di Tommaso et al., 2017), Parsl (Babuji et al., 2019), Pegasus (Deelman et al., 2015), Swift (Wilde et al., 2009)).
-WMSs can be differentiated by how dependency graphs are defined (Pauloski et al., 2024): static configurations files, such as CWL (Crusoe et al., 2022), XML, or YAML; general purpose languages (GPLs); domain specific languages (DSLs); or procedurally through the dynamic execution of a program.
+Parallel computing libraries, such as Dask [53] and Ray [54], provide mechanisms for executing functions in parallel across local resources or distributed systems.
+Similarly, workflow management systems (WMSs) can execute tasks in parallel but also provide mechanisms for defining, optimizing, and monitoring DAG execution (e.g., Airflow [55], Fireworks [56], Makeflow [57], Nextflow [58], Parsl [32], Pegasus [59], Swift [60]).
+WMSs can be differentiated by whether dependency graphs are defined [61] with static configurations files, such as CWL [62], XML, or YAML; general purpose languages (GPLs); domain specific languages (DSLs); or procedurally through the dynamic execution of a program.
 The class of workflows supported by these frameworks have two key limitations that we address: tasks are assumed to be pure (i.e., no side-effects) and programs are static, i.e., they cannot adapt to changing environments over time.
-Actors are computational entities that enable concurrent computing through message passing (Hewitt et al., 1973).
+Actors are computational entities that enable concurrent computing through message passing [20].
 In response to a message, an actor can alter its local state, send messages to other actors, and create new actors.
-No global state means locks and other synchronization primitives are not required.
-Actors can enable stateful computations within traditionally stateless programming models, and are supported in parallel computing frameworks (e.g., Akka (Lightbend, 2009), Dask, Orleans (Bernstein et al., 2014), Ray) and function-as-a-service (FaaS) platforms (e.g., Abaco (Garcia et al., 2020), Azure Service Fabric (Microsoft, 2017), PraaS (Copik et al., 2022)).
-Actor models have been investigated as alternatives for designing computational workflows where communication and coordination are decoupled (Bowers and Ludäscher, 2005).
+No global state means locks and synchronization primitives are not required.
+Actors can enable stateful computations within traditionally stateless programming models, and are supported in parallel computing frameworks (e.g., Akka [63], Dask, Orleans [64], Ray) and function-as-a-service (FaaS) platforms (e.g., Abaco [65], Azure Service Fabric [66], PraaS [67]).
+Actor models have been investigated as alternatives for designing computational workflows where communication and coordination are decoupled [68].
 Our system extends the actor model to support autonomous behaviors and federated deployments.
 Multi-agent systems can enhance or automate scientific processes.
-Early work investigated cooperative agent environments for distributed problem solving with minimal human intervention (Drashansky et al., 1995, 1999).
-Recent work focuses on improving the reasoning capabilities of LLM-backed agents through ontological knowledge graphs and multi-agent systems (Ghafarollahi and Buehler, 2024) and tool-augmented LLMs (Ma et al., 2024).
-Increasingly popular is the use of multi-agent conversations, in which multiple role-specialized agents interact to collaborate, coordinate, or compete towards goals (Wu et al., 2023).
-These systems enhance LLM-based tools through better reasoning (Du et al., 2023), validation (Wu et al., 2024), and divergent thinking (Liang et al., 2024), prompting rapid development of frameworks such as LangGraph (LangChain, 2024), Microsoft AutoGen (Wu et al., 2023), OpenAI Swarm (OpenAI, 2024), and Pydantic Agents (Pydantic, 2024).
+Early work investigated cooperative agent environments for distributed problem solving with minimal human intervention [69, 70].
+Recent work focuses on improving the reasoning capabilities of LLM-backed agents through ontological knowledge graphs and multi-agent systems [71] and tool-augmented LLMs [72].
+Increasingly popular is the use of multi-agent conversations, in which multiple role-specialized agents interact to collaborate, coordinate, or compete towards goals [11].
+These systems enhance LLM-based tools through better reasoning [73], validation [74], and divergent thinking [75], prompting rapid development of frameworks such as LangGraph [12], Microsoft AutoGen [11], OpenAI Swarm [13], and Pydantic Agents [76].
 Subsequently, interest in standardizing agent protocols has developed.
-Anthropic’s Model Context Protocol (MCP) (Anthropic, 2024) defines structured interaction between humans/tools and AI models.
-Google’s Agent2Agent (A2A) Protocol (Google, 2025) focuses on structured interaction between autonomous agents; each agent serves an HTTP endpoint which is impracticable for many scientific workflows.
-Multi-agent conversations can proxy scientists in iterative scientific processes—brainstorming ideas, planning experiments, and reasoning about results (Boiko et al., 2023; Wang et al., 2023; Gao et al., 2024; Google Research, 2025)—but these aforementioned systems are designed for local or cloud-native applications and lack the features necessary to deploy agents across federated research infrastructure.
+Anthropic’s Model Context Protocol (MCP) [77] defines structured interaction between humans/tools and AI models.
+Google’s Agent2Agent (A2A) Protocol [78] focuses on structured interaction between autonomous agents; each agent serves an HTTP endpoint which is impracticable for many scientific workflows.
+Multi-agent conversations can proxy scientists in iterative scientific processes—brainstorming ideas, planning experiments, and reasoning about results [79, 9, 10, 80]—but these aforementioned systems are designed for local or cloud-native applications and lack the features necessary to deploy agents across federated research infrastructure.
 We focus on the systems-level challenges of representing and deploying diverse agent types and agentic workflows across heterogeneous environments rather than the applied use of LLMs for workflow steering.
-7. Conclusion & Future Work
+VIII Conclusion & Future Work
 Advancements in AI, coupled with concurrent advancements in self-driving laboratories, high performance computing, and research data management, open the door for truly autonomous scientific discovery.
@@
 We presented solutions to three key challenges: representing and programming agents; communicating among agents; and executing agents across diverse resources.
-Our evaluations demonstrate that Academy can support high-performance workflows, and three case studies highlight the advantages of agentic workflow design.
-In future work, we will explore scoped authentication to control which agents can invoke others, enabling the creation of agent marketplaces where access can be granted, revoked, or delegated.
-We also plan to expand agent discovery with additional metadata to support AI-steered workflows in which LLMs autonomously identify and use available agents.
-Recording the relative ordering of agent events (i.e., messages received and state transitions), as in Instant Replay (Leblanc and Mellor-Crummey, 1987), can support provenance within agentic workflows.
-Last but not least, we will work across scientific research communities to assemble agents for different purposes, and with research facilities to identify obstacles to agent use that may motivate further developments in Academy.
-8. Acknowledgments
+Our evaluations demonstrate that Academy can support high-performance workflows, and four case studies highlight the advantages of agentic workflow design.
+IX Acknowledgments
 This research was supported in part by the National Science Foundation under Grants 2004894 and 2209919.